### v1

# Value-Free Policy Optimization via Reward Partitioning

Abstract
Single-trajectory reinforcement learning (RL) methods aim to optimize policies from datasets consisting of (prompt, response, reward) triplets, where scalar rewards are directly available. This supervision format is highly practical, as it mirrors real-world human feedback, such as thumbs-up/down signals, and avoids the need for structured preference annotations. In contrast, pairwise preference-based methods like Direct Preference Optimization (DPO) rely on datasets with both preferred and dispreferred responses, which are harder to construct and less natural to collect. Among single-trajectory approaches, Direct Reward Optimization (DRO) has shown strong empirical performance due to its simplicity and stability. However, DRO requires approximating a value function, which introduces several limitations: high off-policy variance, coupling between policy and value learning, and a lack of absolute supervision on the policy itself. We introduce Reward Partitioning Optimization (RPO), a new method that resolves these limitations by removing the need to model the value function. Instead, RPO normalizes observed rewards using a partitioning approach estimated directly from data. This leads to a straightforward supervised learning objective on the policy, with no auxiliary models and no joint optimization. RPO provides direct and stable supervision on the policy, making it robust and easy to implement in practice.
We validate RPO on scalar-feedback language modeling tasks using Flan-T5 encoder-decoder models. Our results demonstrate that RPO outperforms existing single-trajectory baselines such as DRO and Kahneman-Tversky Optimization (KTO). These findings confirm that RPO is a simple, effective, and theoretically grounded method for single-trajectory policy optimization. Our code is publicly available at https://github.com/b-faye/rpo.
I Introduction
The dominant approach for aligning large language models (LLMs), whether through reinforcement learning from human feedback (RLHF) [1, 2, 3] or direct preference optimization, is to learn from preference or reward data [4, 5, 6, 7]. In RLHF, models are first trained via supervised fine-tuning (SFT) [8], followed by learning a reward model (RM) from human-labeled preference pairs (e.g., two responses to the same prompt ranked by quality), and finally aligned with reinforcement learning (e.g., PPO [9], TRPO [10]) to maximize the RM’s score. This pipeline is central to the training of models such as InstructGPT [11], GPT-4 [12], and Claude 3 [13].
While effective, RLHF presents key limitations: collecting preference data and training reward models is costly, and the RL step is sample-inefficient and brittle. To reduce reliance on human annotation, recent work has proposed Reinforcement Learning from AI Feedback (RLAIF), where feedback is generated by models rather than humans. RLAIF includes strategies such as distilling AI feedback into a reward model [14], using an LLM directly as a reward function [15, 16], or leveraging self-evaluation [17, 18]. These techniques offer scalability but still depend on large models as evaluators, reintroducing cost and inference complexity.
A parallel line of work avoids reward models and RL altogether, by directly optimizing policies from preference data via supervised learning. These methods—such as DPO [4], GPO [5], and SLiC [19]—contrast response pairs to a prompt and define a preference-based loss. This strategy enables simple and stable training, and has been adopted in models like LLaMA 3 [20], Qwen 2 [21], and Nemotron-4 [22].
However, collecting high-quality pairwise preference data is increasingly difficult. As LLMs improve, distinguishing between good responses becomes subtler, demanding more qualified annotators and driving up labeling costs. This makes scaling preference-based methods impractical at the frontier.
To address this, recent work proposes single-trajectory learning, where data consists of a prompt, a response, and a scalar reward (e.g., thumbs-up/down). This format is simpler and more natural for annotation. Methods such as KTO [7] and DRO [6] leverage this setup to train policies directly from reward-labeled completions.
Among them, DRO has shown strong empirical results by regressing the optimality condition from KL-regularized RL. However, DRO suffers from several limitations: it requires learning a value function from off-policy data, the joint optimization over policy and value is unstable, and the policy is only supervised relatively to a reference model, not directly by reward.
We introduce Reward Partitioning Optimization (RPO), a new method for single-trajectory policy optimization. RPO eliminates the need to estimate a value function by computing an empirical normalization over observed rewards. This enables a direct supervised loss on the policy, using only prompt-response-reward triplets and a reference model.
RPO is simple to implement, avoids instability from value approximation, and provides absolute reward supervision. It is fully offline-compatible and scalable to large datasets. Our contributions are:
  • We analyze theoretical and practical limitations of DRO for single-trajectory learning.
  • We propose RPO, a new method that replaces value estimation with empirical reward normalization.
  • We provide a formal derivation showing RPO approximates the optimal policy structure.
  • We validate RPO on instruction-tuning tasks using T5 encoder-decoder models [23], demonstrating strong performance against DRO and KTO baselines.
II Background
In this section, we review existing paradigms for aligning large language models (LLMs) with human feedback, focusing on three main categories: reinforcement learning from human feedback (RLHF), direct preference optimization (DPO-style), and direct reward optimization (DRO-style). We examine their assumptions, limitations, and motivation for a more general framework that can directly leverage scalar reward data.
II-A Reinforcement Learning from Human Feedback (RLHF)
RLHF has become the dominant paradigm for aligning large language models (LLMs) with human preferences. It proceeds in three main stages:
  • Supervised Fine-Tuning (SFT): A reference policy $\pi_{\text{ref}}$ is trained using standard supervised learning to mimic high-quality human-written responses.
  • Reward Model (RM) Training: Given a dataset $\mathcal{D}$ of pairwise preferences $(x,y^{\text{win}},y^{\text{lose}})$—where $x$ is a prompt, and $y^{\text{win}}\succ y^{\text{lose}}$ indicates a preferred response—the goal is to learn a scalar reward function $r_{\phi}(x,y)$ that reflects human preferences. This is typically modeled using a Bradley-Terry [24] formulation:
    $\footnotesize\mathbb{P}[y^{\text{win}}\succ y^{\text{lose}}\mid x]=\sigma(r_{\phi}(x,y^{\text{win}})-r_{\phi}(x,y^{\text{lose}})),$ (1)
    where $\sigma$ is the logistic sigmoid. The reward model is trained by minimizing the negative log-likelihood:
    $\footnotesize\mathcal{L}_{\text{RM}}(r_{\phi})=\mathbb{E}_{(x,y^{\text{win}},y^{\text{lose}})\sim\mathcal{D}}[-\log\sigma(r_{\phi}(x,y^{\text{win}})-r_{\phi}(x,y^{\text{lose}}))].\$ (2)
  • Policy Optimization: A policy $\pi_{\theta}$ is then optimized to generate outputs that maximize the learned reward $r_{\phi}(x,y)$ while staying close to the reference policy $\pi_{\text{ref}}$. This is formalized as the following Kullback-Leibler (KL)-regularized objective:
    $\footnotesize\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},\,y\sim\pi_{\theta}(\cdot|x)}\left[r_{\phi}(x,y)-\tau\cdot\text{KL}(\pi_{\theta}(y|x)\parallel\pi_{\text{ref}}(y|x))\right],$ (3)
    where $\tau>0$ is a regularization coefficient that balances reward maximization and policy deviation. Since this objective is not differentiable with respect to $\pi_{\theta}$, reinforcement learning algorithms such as Proximal Policy Optimization (PPO) [9] are used to perform the optimization.
Despite its success, RLHF is resource-intensive and challenging to implement. Training the reward model demands significant human supervision and is susceptible to out-of-distribution generalization errors that can mislead the policy. Furthermore, reinforcement learning algorithms add substantial computational overhead and are sensitive to hyperparameter tuning. These drawbacks have spurred interest in alternative alignment strategies that bypass explicit reinforcement learning.
II-B Direct Preference Optimization (DPO-style)
To address the high cost and complexity of full reinforcement learning, a family of recent approaches—such as DPO [4], IPO [25], and SLiC [19]—propose optimizing policies directly from preference data without learning a separate reward model or relying on reinforcement learning algorithms like PPO.
Given a dataset of pairwise preferences $\mathcal{D}=\{(x,y^{\text{win}},y^{\text{lose}})\}_{i=1}^{N}$, where $x$ is a prompt and $y^{\text{win}}\succ y^{\text{lose}}$ indicates a preferred completion, these methods define an objective that compares the log-likelihoods of the winning and losing responses under the learned policy $\pi_{\theta}$, relative to a fixed reference policy $\pi_{\text{ref}}$. The general form of the loss is:
$\footnotesize\mathcal{L}_{\text{pref}}(\theta)=\mathbb{E}_{(x,y^{\text{win}},y^{\text{lose}})\sim\mathcal{D}}\left[f\left(\beta\left(\log\frac{\pi_{\theta}(y^{\text{win}}|x)}{\pi_{\text{ref}}(y^{\text{win}}|x)}-\log\frac{\pi_{\theta}(y^{\text{lose}}|x)}{\pi_{\text{ref}}(y^{\text{lose}}|x)}\right)\right)\right],$ (4)
where $\beta>0$ is a temperature parameter and $f$ is a shaping function that varies by method:
  • DPO uses the logistic loss: $f(z)=\log(1+e^{-z})$;
  • SLiC uses a hinge loss: $f(z)=\max(0,1-z)$;
  • IPO uses a squared loss centered at $z=1$: $f(z)=(z-1)^{2}$.
This loss encourages the policy $\pi_{\theta}$ to assign higher probability to preferred responses than to dispreferred ones, while penalizing large deviations from the reference policy $\pi_{\text{ref}}$. In essence, it minimizes a divergence between the preference-consistent ranking and the policy-induced ranking, thereby aligning the model’s behavior with human preferences without requiring a learned reward function or reinforcement learning procedure.
These methods can be interpreted as KL-regularized preference optimization, where the model is trained to favor preferred completions relative to the reference while remaining close to $\pi_{\text{ref}}$. Crucially, they bypass the need for an explicit reward model and reinforcement learning, leading to simpler and more stable training.
However, their reliance on manually curated preference datasets poses a practical limitation. In real-world deployment, feedback often comes in the form of scalar or binary reward signals, i.e., as $(x,y,r)$ tuples rather than preferences—motivating alternative approaches capable of leveraging such logged data directly.
II-C Direct Reward Optimization (DRO-style)
DRO-style refers to a class of methods that directly optimize a policy from scalar rewards rather than pairwise preferences. Given a dataset $\mathcal{D}=\{(x,y,r)\}_{i=1}^{N}$, where $x$ is a prompt, $y$ a model response, and $r\in\{-1,+1\}$ a scalar reward indicating human feedback, DRO-style methods aim to maximize the utility of desirable responses while minimizing that of undesirable ones—without relying on an explicit reward model or reinforcement learning algorithm.
A representative example of this approach is the KTO [7] method, which draws inspiration from prospect theory [26] to model human preferences with loss aversion and risk sensitivity.
Let $\pi_{\theta}$ denote the current policy and $\pi_{\mathrm{ref}}$ a fixed reference policy. Define the log-likelihood ratio between the two as:
$\footnotesize r_{\theta}(x,y)=\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}.$ (5)
Next, define a reference point $z_{0}$ as the expected log-likelihood ratio over sampled responses:
$\footnotesize z_{0}=\mathbb{E}_{y^{\prime}\sim\pi_{\theta}(\cdot\mid x)}\left[\log\frac{\pi_{\theta}(y^{\prime}\mid x)}{\pi_{\mathrm{ref}}(y^{\prime}\mid x)}\right]=\mathrm{KL}(\pi_{\theta}(\cdot\mid x)\|\pi_{\mathrm{ref}}(\cdot\mid x)).$ (6)
To model asymmetric treatment of gains and losses, KTO introduces the following components:
  • A temperature parameter $\beta>0$ to control the sharpness of value sensitivity;
  • A logistic function $\sigma(z)=\frac{1}{1+e^{-z}}$ for smooth transition between reward regions;
  • Distinct loss aversion coefficients $\lambda_{D}$ (for desirable responses) and $\lambda_{U}$ (for undesirable ones).
The value function $V(x,y)$ is then defined as:
$\footnotesize V(x,y)=\begin{cases}\lambda_{D}\cdot\sigma\left(\beta\left[r_{\theta}(x,y)-z_{0}\right]\right),&\text{if }r=+1,\\ \lambda_{U}\cdot\sigma\left(\beta\left[z_{0}-r_{\theta}(x,y)\right]\right),&\text{if }r=-1.\end{cases}$ (7)
Finally, the KTO objective minimizes the following expected loss over the data distribution:
$\footnotesize\mathcal{L}_{\mathrm{KTO}}(\theta)=\mathbb{E}_{(x,y,r)\sim\mathcal{D}}\left[\lambda_{r}-v(x,y)\right],$ (8)
where $\lambda_{r}=\lambda_{D}$ if $r=+1$ and $\lambda_{U}$ if $r=-1$.
This formulation allows KTO to smoothly reward desirable generations and penalize undesirable ones, while incorporating human-like asymmetries in preference treatment. Unlike binary cross-entropy or advantage-weighted likelihood objectives, KTO provides a principled, prospect-theoretic loss tailored to scalar reward data.
KTO models human risk aversion but relies on strong assumptions and optimizes only the policy $\pi_{\theta}$, ignoring a separate value function $V(x)$. In contrast, DRO [6] jointly optimizes both $\pi$ and $V$, offering a more principled and general framework.
Given a dataset $\mathcal{D}=\{(x,y,r)\}_{i=1}^{N}$ where $r$ is a scalar reward, DRO formulates policy learning as a KL-regularized reward maximization problem:
$\footnotesize\pi^{*}=\arg\max_{\pi}\;\mathbb{E}_{x\sim\rho,y\sim\pi(\cdot|x)}\left[r(x,y)-\tau\cdot\mathrm{KL}\left(\pi(\cdot|x)\|\pi_{\mathrm{ref}}(\cdot|x)\right)\right],$ (9)
where $\pi_{\mathrm{ref}}$ is a fixed reference policy and $\tau>0$ is a temperature hyperparameter controlling the strength of regularization.
The optimal policy $\pi^{*}$ can be expressed in closed form as:
$\footnotesize\pi^{*}(y|x)=\pi_{\mathrm{ref}}(y|x)\cdot\exp\left(\frac{1}{\tau}r(x,y)-\frac{1}{\tau}V^{*}(x)\right),$ (10)
where $V^{*}(x)$ is the soft value function, defined by
$\footnotesize V^{*}(x)=\tau\log\mathbb{E}_{y\sim\pi_{\mathrm{ref}}(\cdot|x)}\left[\exp\left(\frac{1}{\tau}r(x,y)\right)\right].$ (11)
Rearranging, this identity links reward, value, and policy:
$\footnotesize r(x,y)=V^{*}(x)+\tau\cdot\log\frac{\pi^{*}(y|x)}{\pi_{\mathrm{ref}}(y|x)}.$ (12)
To approximate the optimal $(\pi^{*},V^{*})$, DRO minimizes the squared Bellman residual:
$\footnotesize\mathcal{L}_{\mathrm{DRO}}(\theta,\phi)=\frac{1}{2}\mathbb{E}_{(x,y,r)\sim\mathcal{D}}\left[\left(r-V_{\phi}(x)-\tau\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}\right)^{2}\right],$ (13)
where $\pi_{\theta}$ and $V_{\phi}$ denote parameterized policy and value functions respectively.
In practice, this loss is optimized jointly over parameters $\theta$ and $\phi$ using the observed data samples $\{(x_{i},y_{i},r_{i})\}_{i=1}^{n}$ as:
^$\footnotesize\hat{\mathcal{L}}_{\mathrm{DRO}}(\theta,\phi)=\frac{1}{2N}\sum_{i=1}^{N}\left(r_{i}-V_{\phi}(x_{i})-\tau\log\frac{\pi_{\theta}(y_{i}|x_{i})}{\pi_{\mathrm{ref}}(y_{i}|x_{i})}\right)^{2}.$ (14)
This principled framework optimizes policies directly from scalar rewards by jointly learning a value function and applying KL regularization for stability. However, this joint optimization often suffers from instability, especially when the data distribution differs from the policy, as value estimation errors can bias training. To address these issues, we propose RPO, which retains DRO’s strengths but removes the need to estimate a value function, simplifying training and enhancing robustness.
III Method
Direct policy optimization in the RLHF setting with KL regularization commonly relies on the classical exponential form of the optimal policy:
$\footnotesize\pi^{*}(y|x)=\frac{\pi_{\mathrm{ref}}(y|x)\exp\left(\frac{1}{\tau}r(x,y)\right)}{Z(x)},$ (15)
where the partition function $Z(x)$ is
$\footnotesize Z(x)=\sum_{y^{\prime}}\pi_{\mathrm{ref}}(y^{\prime}|x)\exp\left(\frac{1}{\tau}r(x,y^{\prime})\right).$ (16)
Current methods such as DRO learn both a policy $\pi_{\theta}$ and a value function $V_{\phi}$ parameterized by $\phi$, where the soft value function relates to the partition function as $V^{*}(x)=\tau\log Z(x)$. However, this joint learning introduces several difficulties: the value function $V_{\phi}$ can be unstable to approximate, the parameters $(\theta,\phi)$ are tightly coupled, and optimization becomes more complex.
To address these challenges, we propose an alternative method, Reward Partitioning Optimization (RPO), which completely removes the need for an explicit value function model. Instead, RPO performs direct regression on the log-ratio of policies by empirically estimating the partition function, leveraging its analytical form for simpler and more stable training.
Starting from the optimality condition of the policy:
$\footnotesize r(x,y)-V^{*}(x)=\tau\log\frac{\pi^{*}(y|x)}{\pi_{\mathrm{ref}}(y|x)},$ (17)
where the value function $V^{*}(x)$ normalizes the exponentiated rewards, expressible exactly as:
$\footnotesize V^{*}(x)=\tau\log\sum_{y^{\prime}}\pi_{\mathrm{ref}}(y^{\prime}|x)\exp\left(\frac{1}{\tau}r(x,y^{\prime})\right).$ (18)
Instead of learning $V^{*}$ via a separate neural network, RPO estimates it empirically by aggregating samples with the same prompt $x$. Let $\mathcal{I}_{x}=\{i:x_{i}=x\}$ index all samples for prompt $x$. Then, the empirical partition function and value are:
$\footnotesize\widehat{Z}(x)=\sum_{j\in\mathcal{I}_{x}}\pi_{\mathrm{ref}}(y_{j}|x)\exp\left(\frac{1}{\tau}r(x,y_{j})\right),\quad\widehat{V}(x)=\tau\log\widehat{Z}(x).$ (19)
RPO minimizes the squared error between the log-policy ratio and the scaled centered reward, optimizing only over the policy parameters $\theta$. Formally, the loss can be expressed as the expectation:
$\footnotesize\mathcal{L}_{\mathrm{RPO}}(\theta)=\frac{1}{2}\mathbb{E}_{(x,y,r)\sim\mathcal{D}}\left[\left(\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}-\frac{1}{\tau}(r(x,y)-V^{*}(x))\right)^{2}\right].$ (20)
In practice, we approximate this with the empirical sum over the training set $\{(x_{i},y_{i},r_{i})\}_{i=1}^{n}$ and use $\widehat{V}(x)$ as an estimate for $V^{*}(x)$:
$\footnotesize\mathcal{L}_{\mathrm{RPO}}(\theta)=\frac{1}{2n}\sum_{i=1}^{n}\left(\log\frac{\pi_{\theta}(y_{i}|x_{i})}{\pi_{\mathrm{ref}}(y_{i}|x_{i})}-\frac{1}{\tau}\left(r_{i}-\widehat{V}(x_{i})\right)\right)^{2}.$ (21)
Defining the residual for sample $i$ as:
$\footnotesize\delta_{i}(\theta):=\log\frac{\pi_{\theta}(y_{i}|x_{i})}{\pi_{\mathrm{ref}}(y_{i}|x_{i})}-\frac{1}{\tau}\left(r_{i}-\widehat{V}(x_{i})\right),$ (22)
we compute the gradient of the loss with respect to the policy logits $z_{\theta}(y|x)$, where the policy is parameterized via a softmax:
$\footnotesize\pi_{\theta}(y|x)=\frac{\exp(z_{\theta}(y|x))}{\sum_{y^{\prime}}\exp(z_{\theta}(y^{\prime}|x))}.$ (23)
The gradient of the loss is then:
$\footnotesize\nabla_{z_{\theta}(y|x)}\mathcal{L}_{\mathrm{RPO}}(\theta)=\frac{1}{n}\sum_{i:x_{i}=x}\delta_{i}(\theta)\cdot\nabla_{z_{\theta}(y|x)}\log\pi_{\theta}(y_{i}|x).$ (24)
Using the standard softmax gradient, we have:
$\footnotesize\nabla_{z_{\theta}(y|x)}\log\pi_{\theta}(y_{i}|x)=\mathbf{1}_{y=y_{i}}-\pi_{\theta}(y|x),$ (25)
where $\mathbf{1}_{y=y_{i}}$ is the indicator function.
Substituting into Equation (24), the final gradient expression becomes:
$\footnotesize\nabla_{z_{\theta}(y|x)}\mathcal{L}_{\mathrm{RPO}}(\theta)=\frac{1}{n}\sum_{i:x_{i}=x}\delta_{i}(\theta)\left(\mathbf{1}_{y=y_{i}}-\pi_{\theta}(y|x)\right).$ (26)
This gradient has an intuitive interpretation as the weighted discrepancy between the one-hot target and the policy prediction, weighted by the residual error $\delta_{i}(\theta)$. Crucially, RPO avoids backpropagation through the value estimate $\widehat{V}(x)$, resulting in a simpler, more stable training procedure that retains the theoretical benefits of DRO without learning an explicit value function.
1
Input: Dataset $\mathcal{D}=\{(x_{i},y_{i},r_{i})\}_{i=1}^{N}$, initial policy $\pi_{\theta}$, reference policy $\pi_{\mathrm{ref}}$, temperature $\tau>0$, number of epochs $T$, learning rate $\eta$
Output: Trained policy $\pi_{\theta}$
2
3Group indices by prompt: $\forall x,\mathcal{I}_{x}=\{i:x_{i}=x\}$;
4
5for $t\leftarrow 1$ to $T$ do
6       foreach batch $B$ from $\mathcal{D}$ do
7             foreach $(x_{i},y_{i},r_{i})\in B$ do
8                   Compute $\ell_{i}\leftarrow\log\pi_{\theta}(y_{i}|x_{i})$;
9                   Compute $\ell_{i}^{\mathrm{ref}}\leftarrow\log\pi_{\mathrm{ref}}(y_{i}|x_{i})$;
10
11                  Estimate empirical partition: $\widehat{Z}(x_{i})=\sum_{j\in\mathcal{I}_{x_{i}}}\pi_{\mathrm{ref}}(y_{j}|x_{i})\exp\left(\frac{1}{\tau}r(x_{i},y_{j})\right)$;
12
13                  Compute $\widehat{V}(x_{i})\leftarrow\tau\log\widehat{Z}(x_{i})$;
14                   Compute $t_{i}\leftarrow\frac{r_{i}-\widehat{V}(x_{i})}{\tau}$;
15                   Compute $\ell_{\mathrm{RPO},i}\leftarrow\frac{1}{2}\left(\ell_{i}-\ell_{i}^{\mathrm{ref}}-t_{i}\right)^{2}$;
16
17            Update $\theta$ by gradient descent on $\frac{1}{|B|}\sum_{i\in B}\ell_{\mathrm{RPO},i}$;
18
19
20return $\pi_{\theta}$;
Algorithm 1 RPO Optimization
To operationalize the RPO training procedure described above, we present a concrete algorithmic implementation (Algorithm 1). The algorithm iteratively estimates the target log-ratio using a partition function approximation based on the reference policy, computes the residual loss for each sample, and updates the policy parameters via gradient descent.
Property DRO RPO
Explicit $V$ estimation Yes (model $V_{\phi}$) No (empirical estimation)
$\pi$$V$ coupling Strong Weak
Optimization complexity Medium to high Simpler, more stable
Robustness to estimation errors Sensitive Less sensitive
Practical deployment Requires tuning More direct, no $V$ tuning
TABLE I: Comparison between DRO and RPO methods.
To further highlight the distinctions between the proposed RPO method and standard distributionally robust optimization (DRO) approaches, Table I summarizes key differences in terms of implementation, coupling, and practical considerations.
The RPO method formulates policy learning as a direct regression on the log-ratio between target and reference policies, using rewards and an empirical partition estimation. This removes the need to train a value function, simplifying the learning procedure and enhancing numerical stability.
IV Experiments
IV-A Dataset and Base Models
[…]
We adopt the Flan-T5 model family [28], a series of encoder-decoder transformers fine-tuned for instruction following. To examine the effect of model scale on RPO performance, we experiment with three variants: Flan-T5 Small (77M parameters), Flan-T5 Large (783M parameters), and Flan-T5 XL (2.85B parameters). For clarity, we refer to these models as Small, Large, and XL, respectively, throughout the remainder of the paper.
All models are initialized from instruction-tuned checkpoints following the FLAN recipe [29], serving as the reference policy $\pi_{\text{ref}}$. This consistent initialization allows for direct and fair comparison of RPO across model sizes under the same training framework.
IV-B Evaluation
To assess model performance, we construct a validation set of 1,000 examples from the UltraFeedback dataset [27]. These examples are selected based on their high reward scores (equal to 10.0), which indicates strong user preference and high-quality completions. This curated subset serves as the basis for two complementary evaluation methods.
Automatic Metric-Based Validation. The first evaluation strategy is conducted during training to monitor model progression. We periodically evaluate the model outputs using standard automatic metrics, including BLEU, ROUGE-L, METEOR, BERTScore, SacreBLEU, and Toxicity. These metrics—commonly used in text generation—provide surface-level assessments of similarity and relevance to reference completions. While useful for tracking learning dynamics and verifying general quality trends, such metrics are limited in capturing deeper alignment, preference satisfaction, or nuanced helpfulness.
LLM-Based Preference Evaluation. To address the limitations of automatic metrics in the reinforcement learning with human feedback (RLHF) setting, we adopt a second evaluation method based on pairwise comparisons judged by large language models (LLMs). This type of evaluation, known as LLM-as-a-judge [30], provides a more human-aligned assessment of response quality. After training, we compare model completions against a reference using state-of-the-art LLMs—Mistral v3 [31] and LLaMA v3 [32]—as impartial judges.
Each evaluation instance presents the judge with an instruction and two responses (A and B), and the model is asked to select the more helpful and relevant one. The evaluation prompt used for this pairwise judgment is as follows:
You are an impartial assistant that judges responses based on helpfulness and relevance.
You are provided an instruction and two responses, A and B.
Decide which response better fulfills the instruction in terms of helpfulness and relevance.
Answer strictly with "A" or "B" only. Do NOT provide explanations or other text.
Instruction:
{instruction}
Response A:
{response_a}
Response B:
{response_b}
Your answer (A or B): This method allows for a fine-grained, preference-sensitive evaluation that better reflects the goals of RLHF-based optimization. It enables us to assess whether models trained with RPO produce responses that are genuinely preferred by LLM judges, rather than merely similar to reference outputs.
IV-C Training Setup and Baseline Comparison
We train all models using the AdamW [33] optimizer with a learning rate of $1\times 10^{-4}$. The learning rate scheduler follows a linear warmup and decay schedule, with 150 warmup steps and total training steps computed as the product of steps per epoch and the number of epochs:
[…]
The scheduler is configured using the get_scheduler function from the Hugging Face Transformers library with linear scheduling.
For evaluating the hyperparameters of our proposed method RPO, we use KTO and DRO as baselines. The temperature parameter $\tau$ is set to 1.0 in all experiments to minimize its impact and focus on intrinsic method differences. Setting $\tau=1.0$ ensures a fair comparison by controlling the sharpness of the policy distribution uniformly.
Figure 1 compares the average training durations in hours across three FLAN-T5 model sizes (Small, Large, XL) for KTO, DRO, and RPO.
Figure 1: Average training time in hours for KTO, DRO, and RPO across FLAN-T5 model sizes: Small (77M parameters), Large (783M parameters), and XL (2.85B parameters). RPO consistently converges faster, demonstrating superior training efficiency.
As illustrated, RPO converges faster than both KTO and DRO for all model scales, highlighting its efficiency and suitability for scalable single-trajectory policy optimization.
IV-D Automatic Metric-Based Validation
During training, we perform automatic validation at various checkpoints using standard natural language generation metrics to evaluate the quality of the generated responses. This validation protocol provides quantitative feedback on the model’s performance in terms of semantic similarity, fluency, and toxicity. We compare three training methods: KTO, DRO, and our proposed method RPO across three model sizes: Small, Large, and XL.
The metrics used include BERTScore F1, ROUGE-L, BLEU, SacreBLEU, METEOR, and average toxicity. Arrows indicate the desired direction for each metric: $\uparrow$ means higher values are better, while $\downarrow$ means lower values are better.
Metric KTO DRO RPO
Small Large XL Small Large XL Small Large XL
BERTScoreF1 0.8499 0.8401 0.7883 0.8562 0.8591 0.8537 0.8661 0.8706 0.8702
ROUGE-L 0.2369 0.2491 0.2480 0.2522 0.3034 0.2502 0.2885 0.3162 0.2969
BLEU 0.0293 0.0438 0.0108 0.0173 0.0274 0.0419 0.0360 0.0464 0.0511
SacreBLEU 2.8740 4.3785 1.3653 1.7218 2.8589 4.2672 3.6013 4.7903 5.1063
METEOR 0.2083 0.2152 0.203 0.2029 0.2337 0.2153 0.2418 0.2632 0.2567
Toxicityavg 0.0133 0.0241 0.0242 0.0146 0.0130 0.0122 0.0125 0.0093 0.0078
TABLE II: Automatic metric-based validation results for KTO, DRO, and our proposed RPO method across different model sizes. Metrics with $\uparrow$ indicate that higher values are better, while $\downarrow$ indicates lower values are preferable. RPO consistently achieves better scores in most metrics, highlighting its effectiveness.
Table II summarizes the validation metrics across different training methods and model scales. We observe that RPO generally outperforms both KTO and DRO in semantic similarity and fluency metrics while maintaining lower toxicity levels. This suggests that RPO enhances the preference alignment of generated responses more effectively than baseline approaches.
However, these standard metrics have inherent limitations in the context of RLHF tasks. Metrics like BLEU and ROUGE focus on n-gram overlap and may fail to capture deeper semantic relevance or instruction adherence. Toxicity scores, while useful, do not fully represent nuanced safety concerns. Consequently, relying solely on these metrics can be misleading when optimizing for user preferences.
To address these limitations, we complement this validation with human-like preference evaluations conducted by large language model (LLM) judges, as described in the following subsection.
IV-E LLM-Based Preference Evaluation
To better evaluate alignment with human preferences, we use two strong LLM-based judges—Mistral and LLaMA—to assess outputs from our proposed RPO method compared to three baselines: SFT, KTO, and DRO. We report the percentage of generations where RPO is preferred, along with standard deviation, across three model scales: Small, Large, and XL.
Model Size SFT KTO DRO
Small 82.78 $\pm$ 1.93 82.60 $\pm$ 2.39 83.97 $\pm$ 2.34
Large 84.20 $\pm$ 1.64 86.53 $\pm$ 1.77 84.86 $\pm$ 2.05
XL 83.01 $\pm$ 2.39 93.62 $\pm$ 5.88 83.19 $\pm$ 0.88
(a) Judgment by Mistral
Model Size SFT KTO DRO
Small 63.66 $\pm$ 1.62 65.76 $\pm$ 0.59 65.26 $\pm$ 0.41
Large 73.02 $\pm$ 0.52 72.30 $\pm$ 1.72 74.81 $\pm$ 0.63
XL 68.43 $\pm$ 1.17 61.18 $\pm$ 9.33 67.15 $\pm$ 3.60
(b) Judgment by LLaMA
TABLE III: Winrate (%) of RPO over baseline methods (SFT, KTO, DRO) using LLM judges. Higher values mean stronger preference for RPO.
Analysis. Table III shows that RPO achieves consistently higher winrates than all baselines across sizes and judges. Under the more discriminative Mistral judge, RPO achieves especially strong margins, often exceeding 80% preference.
Why RPO Wins. KTO lacks explicit value modeling and learns from ranking alone, which limits its generalization. DRO introduces value estimation via a dual-network setup (value + policy), but this increases training complexity and potential instability. In contrast, RPO directly optimizes policy through preference ranking, avoiding the need to model or estimate value explicitly. This simplicity leads to more stable training and stronger performance across scales without additional modeling overhead.
IV-F Impact of Regularisation Parameter $\tau$
The temperature parameter $\tau$ in Algorithm 1 plays a key role in shaping the softness of the implicit value estimation $\widehat{V}(x)$, acting as a regularisation factor between reward signals and policy logits. Lower values of $\tau$ emphasize sharper preferences (more deterministic), while higher values smooth the learned policy’s alignment with reference preferences.
We ablate the impact of $\tau$ by evaluating RPO against SFT for $\tau\in\{0.1,1.0,5.0\}$ using LLM judges (Mistral and LLaMA). The winrate of RPO over SFT is reported in Table IV. Results at $\tau=1.0$ correspond to those in Section IV-E.
Model Size $\boldsymbol{\tau=0.1}$ $\boldsymbol{\tau=1.0}$ $\boldsymbol{\tau=5.0}$
Small 84.45 $\pm$ 1.14 82.78 $\pm$ 1.93 83.03 $\pm$ 1.55
Large 59.50 $\pm$ 19.62 84.20 $\pm$ 1.64 83.72 $\pm$ 1.05
XL 90.24 $\pm$ 6.66 83.01 $\pm$ 2.39 82.81 $\pm$ 3.76
(a) Judgment by Mistral
Model Size $\boldsymbol{\tau=0.1}$ $\boldsymbol{\tau=1.0}$ $\boldsymbol{\tau=5.0}$
Small 78.73 $\pm$ 1.07 63.66 $\pm$ 1.62 66.52 $\pm$ 1.55
Large 56.40 $\pm$ 12.21 73.02 $\pm$ 0.52 74.86 $\pm$ 1.63
XL 82.07 $\pm$ 5.30 68.43 $\pm$ 1.17 66.28 $\pm$ 2.48
(b) Judgment by LLaMA
TABLE IV: Winrate (%) of RPO over SFT across different values of the regularisation parameter $\tau$, measured by pairwise preference judgment.
Interpretation. Table IV highlights that RPO is generally robust to $\tau$ variation, particularly at $\tau=1.0$ and $\tau=5.0$. While $\tau=0.1$ performs well at Small and XL scales, it suffers a drop at the Large scale—especially under the Mistral judge—likely due to instability from overly sharp preference weighting.
Moderate values ($\tau=1.0$) yield consistent and strong results, balancing stability and preference sensitivity. Larger $\tau$ values retain strong performance, showing that RPO does not require precise tuning of the temperature. This is a notable advantage compared to methods that tightly couple policy learning with value estimation or reward calibration.
This supports our core insight: RPO avoids direct value modeling, estimating $\widehat{V}(x)$ implicitly via a reference-weighted partition function. It focuses training entirely on policy logits, sidestepping the instability often observed in methods like DRO which rely on dual-network learning (policy + value), and simplifying optimization compared to actor-critic or reward-model-based alternatives.
V Limitations
While RPO offers a simple and scalable alternative to RLHF-style optimization, it is not without limitations.
Empirical Value Estimation. RPO relies on a distributionally robust objective, where the expected value $\widehat{V}(x)$ is computed empirically over the set of candidate responses for each prompt. This estimation assumes sufficient coverage of high-reward outputs under the reference policy $\pi_{\mathrm{ref}}$. If the dataset lacks diverse or high-quality responses, $\widehat{V}(x)$ may be inaccurate, leading to suboptimal updates. In contrast to methods that explicitly train a reward or value model, RPO does not extrapolate beyond observed samples.
Dataset Quality and Coverage. The success of RPO hinges on the quality and diversity of the dataset $\mathcal{D}$. For prompts with limited variation in responses—or when rewards are noisy or sparse—the partition function $\widehat{Z}(x)$ may underestimate the reward landscape. In such settings, the algorithm may favor local modes or fail to discover beneficial deviations from the reference policy.
Dependence on $\pi_{\mathrm{ref}}$. The reference policy serves a dual role: it defines the baseline from which the updated policy diverges, and it shapes the empirical value function. If $\pi_{\mathrm{ref}}$ is poorly aligned with the target reward distribution, the optimization may be biased or less effective.
In summary, RPO’s simplicity comes at the cost of relying heavily on observed data. Addressing these limitations may involve hybridizing with reward modeling, improving data collection strategies, or introducing learned approximations of $\widehat{V}(x)$ to better generalize beyond the dataset.
VI Conclusion and Future Work
In this work, we introduced RPO—a simple, scalable, and reward-driven fine-tuning method that bypasses the need for explicit reward models or reinforcement learning loops. RPO optimizes a robust policy objective that adjusts log-probabilities to match observed preferences while staying close to a reference policy. Our method demonstrates strong empirical performance, outperforming SFT, DRO, and KTO across multiple model sizes.
RPO achieves these results while being significantly simpler to implement and train. Notably, we show that RPO benefits from higher winrates when judged by both Mistral and LLaMA LLMs, indicating strong generalization across distinct evaluators. Furthermore, ablation on the regularization parameter $\tau$ reveals robustness to hyperparameter variation, especially for larger models.
Future Work. There are several promising directions to extend this work. First, we aim to apply RPO to more powerful base models such as GPT-style architectures or frontier LLMs to assess scalability and alignment at higher capacity. Second, we plan to evaluate RPO on broader preference datasets, including real-world human feedback and multi-turn dialogues. Exploring integration with more structured exploration or dynamic reference policies may further improve coverage and robustness. Finally, hybridizing RPO with lightweight value learning or variance reduction techniques could mitigate its current dependency on dataset quality without sacrificing simplicity.

### v2

@@
 To address the limitations of automatic metrics in the reinforcement learning with human feedback (RLHF) setting, we adopt a second evaluation method based on pairwise comparisons judged by large language models (LLMs).
-This type of evaluation, known as LLM-as-a-judge [30], provides a more human-aligned assessment of response quality.
-After training, we compare model completions against a reference using state-of-the-art LLMs—Mistral v3 [31] and LLaMA v3 [32]—as impartial judges.
-Each evaluation instance presents the judge with an instruction and two responses (A and B), and the model is asked to select the more helpful and relevant one.
-The evaluation prompt used for this pairwise judgment is as follows:
+This approach, commonly referred to as LLM-as-a-judge [30], provides a more human-aligned assessment of response quality.
+After training, we compare model completions using state-of-the-art LLMs—Mistral v3 [31] and LLaMA v3 [32]—as impartial judges.
+Each evaluation instance presents the judge with an instruction and two responses (A and B), and the model is asked to decide which response better fulfills the instruction in terms of helpfulness and relevance.
+If neither response is clearly better—or if both are unhelpful or irrelevant—the model is instructed to choose "Neutral".
+The prompt presented to the LLM judge is as follows:
 You are an impartial assistant that judges responses based on helpfulness and relevance.
-You are provided an instruction and two responses, A and B.
+You are provided with an instruction and two responses, A and B.
 Decide which response better fulfills the instruction in terms of helpfulness and relevance.
-Answer strictly with "A" or "B" only. Do NOT provide explanations or other text.
+If both responses are equally good or equally poor, or if neither is clearly better, answer "Neutral".
+Answer strictly with "A", "B", or "Neutral". Do NOT provide explanations or any additional text.
 Instruction:
@@
 {response_b}
-Your answer (A or B): This method allows for a fine-grained, preference-sensitive evaluation that better reflects the goals of RLHF-based optimization.
-It enables us to assess whether models trained with RPO produce responses that are genuinely preferred by LLM judges, rather than merely similar to reference outputs.
+Your answer (A, B, or Neutral):
+This evaluation setup allows for more robust and preference-sensitive comparisons, better reflecting human feedback and alignment goals in RLHF settings.
+By explicitly permitting neutral responses, we reduce bias from forced binary choices and better capture nuanced or inconclusive cases.
 IV-C Training Setup and Baseline Comparison
@@
 (b) Judgment by LLaMA
-TABLE III: Winrate (%) of RPO over baseline methods (SFT, KTO, DRO) using LLM judges. Higher values mean stronger preference for RPO.
+TABLE III: Winrate (%) of RPO over baseline methods (SFT, KTO, DRO) using LLM judges. Higher values indicate stronger preference for RPO completions.
 Analysis.
@@
 (b) Judgment by LLaMA
-TABLE IV: Winrate (%) of RPO over SFT across different values of the regularisation parameter $\tau$, measured by pairwise preference judgment.
+TABLE IV: Winrate (%) of RPO over SFT across different values of the regularization parameter $\tau$, measured by pairwise preference judgment.
 Interpretation.
@@
 It focuses training entirely on policy logits, sidestepping the instability often observed in methods like DRO which rely on dual-network learning (policy + value), and simplifying optimization compared to actor-critic or reward-model-based alternatives.
-V Limitations
+IV-G Ablation Study: Importance of Empirical Normalization
+To assess the role of empirical normalization in Reward Policy Optimization (RPO), we introduce an ablated variant, RPO-noNorm, where the baseline estimate $\widehat{V}(x)$ is removed and the raw reward $r(x,y)$ is used directly.
+This change eliminates advantage-based learning and removes the normalization effect that stabilizes optimization.
+We compare both variants across temperatures $\tau\in\{0.1,1.0,5.0\}$ and model sizes (Small, Large, XL), using both automatic and human evaluations.
+Automatic Evaluation Metrics. Table V presents results across six metrics. We observe:
+  • RPO consistently outperforms RPO-noNorm across all metrics, sizes, and $\tau$ values.
+  • The performance of RPO is stable with respect to temperature, showing its robustness.
+  • RPO-noNorm fails to benefit from temperature scaling and shows consistently lower scores.
+  • Toxicity increases substantially without normalization, suggesting unstable reward learning.
+RPO RPO-noNorm
+$\tau=0.1$ $\tau=1.0$ $\tau=5.0$ $\tau=0.1$ $\tau=1.0$ $\tau=5.0$
+Metric S L XL S L XL S L XL S L XL S L XL S L XL
+BERTScoreF1 0.8648 0.8702 0.8691 0.8661 0.8706 0.8702 0.8653 0.8697 0.8684 0.7734 0.7681 0.7702 0.7741 0.7689 0.7712 0.7726 0.7705 0.7723
+ROUGE-L ↑ 0.2861 0.3143 0.2951 0.2885 0.3162 0.2969 0.2872 0.3150 0.2958 0.2063 0.2028 0.2051 0.2089 0.2043 0.2067 0.2115 0.2078 0.2101
+BLEU ↑ 0.0354 0.0460 0.0503 0.0360 0.0464 0.0511 0.0348 0.0456 0.0498 0.0130 0.0127 0.0139 0.0135 0.0129 0.0143 0.0131 0.0132 0.0137
+SacreBLEU ↑ 3.5421 4.7650 5.0228 3.6013 4.7903 5.1063 3.5470 4.7582 5.0157 1.2360 1.2203 1.2632 1.2743 1.2312 1.2914 1.2634 1.2398 1.2665
+METEOR ↑ 0.2401 0.2619 0.2551 0.2418 0.2632 0.2567 0.2403 0.2626 0.2559 0.1749 0.1694 0.1726 0.1762 0.1708 0.1736 0.1780 0.1719 0.1741
+Toxicityavg 0.0129 0.0095 0.0079 0.0125 0.0093 0.0078 0.0130 0.0097 0.0081 0.0243 0.0250 0.0247 0.0245 0.0251 0.0249 0.0239 0.0244 0.0242
+TABLE V: Automatic metric comparison of RPO and RPO-noNorm across temperatures $\tau\in\{0.1,1.0,5.0\}$ and model sizes.
+S, L, and XL refer to Small, Large, and Extra-Large model variants, respectively.
+RPO consistently achieves stronger performance and lower toxicity across all configurations.
+Temperature $\tau$ has minimal effect compared to normalization.
+These results demonstrate that empirical normalization via $\widehat{V}(x)$ is essential for robust and high-quality optimization.
+Without it, models suffer large drops in fluency, adequacy, and safety—effects that cannot be corrected by tuning $\tau$ alone.
+Human Evaluation via LLM Judges.
+Table VI reports pairwise win-rates of RPO over RPO-noNorm as judged by Mistral and LLaMA.
+RPO is preferred in over 90% of comparisons across all $\tau$ and model sizes.
+This reinforces that empirical normalization not only improves automatic metrics but is also recognized as producing superior outputs by strong language models.
+Model Size $\boldsymbol{\tau=0.1}$ $\boldsymbol{\tau=1.0}$ $\boldsymbol{\tau=5.0}$
+Small 96.22 $\pm$ 0.91 95.04 $\pm$ 0.84 95.77 $\pm$ 1.02
+Large 94.85 $\pm$ 0.77 96.13 $\pm$ 0.66 95.90 $\pm$ 0.58
+XL 97.34 $\pm$ 0.52 95.81 $\pm$ 0.63 96.12 $\pm$ 0.49
+(a) Judgment by Mistral
+Model Size $\boldsymbol{\tau=0.1}$ $\boldsymbol{\tau=1.0}$ $\boldsymbol{\tau=5.0}$
+Small 92.75 $\pm$ 1.13 91.34 $\pm$ 1.25 91.80 $\pm$ 0.97
+Large 90.64 $\pm$ 1.31 91.76 $\pm$ 0.89 92.08 $\pm$ 0.74
+XL 94.21 $\pm$ 1.03 92.99 $\pm$ 0.85 93.17 $\pm$ 0.77
+(b) Judgment by LLaMA
+TABLE VI: LLM preference win-rate (%) of RPO over RPO-noNorm across temperatures and model sizes. RPO is consistently preferred, confirming the necessity of normalization.
+In conclusion, empirical normalization is a decisive factor in the effectiveness of RPO.
+Removing $\widehat{V}(x)$ leads to clear and consistent degradation in both automatic metrics and human preference.
+While temperature $\tau$ modulates exploration, it cannot substitute the stabilizing effect of normalization, making it a critical design choice in reward optimization.
+V Qualitative Study of Prompt Completions
+We conduct a qualitative analysis to illustrate the effects of response optimization strategies on real-world prompts.
+Specifically, we analyze a subset of prompts where both LLaMA and Mistral models demonstrate a preference for completions generated using our RPO method over baseline methods like DRO and KTO.
+In each case, human preferences and reward models agree in ranking RPO higher due to its more complete, helpful, and aligned responses.
+DRO and KTO produce reasonable but weaker completions, often lacking nuance, informativeness, or coherence.
+All examples are generated using the XL variant of each model, as it offers the highest capacity and clearest separation in quality among methods.
+Prompt 1: “How can managers build trust with remote teams while maintaining accountability?”
+RPO: Managers should build trust through consistent communication, transparency about goals, and empathy toward individual circumstances.
+Regular check-ins that focus on support rather than surveillance help establish accountability without eroding trust.
+DRO: Build trust with open communication. Hold people accountable by setting deadlines.
+KTO: Use video calls and check-ins to build trust and keep track of work.
+Prompt 2: “Explain why excessive screen time may negatively affect children’s cognitive development.”
+RPO: Excessive screen time in children is associated with reduced attention span, impaired sleep, and lower engagement in cognitively stimulating activities such as reading or creative play.
+These factors may contribute to slower cognitive and emotional development.
+DRO: Too much screen time can lead to worse focus and bad sleep habits in children.
+KTO: It may reduce focus and sleep.
+Prompt 3: “Is nuclear energy a viable solution for climate change mitigation?”
+RPO: Nuclear energy offers a low-carbon, high-output alternative to fossil fuels and can significantly reduce greenhouse gas emissions.
+While it presents challenges such as waste disposal and high initial costs, it remains a viable component of a broader climate strategy.
+DRO: Nuclear energy is low-carbon but has risks like waste and cost.
+KTO: It helps reduce emissions but is expensive.
+Prompt 4: “What are some effective strategies for learning a new language quickly?”
+RPO: Immersion, daily practice, spaced repetition systems, and conversational exposure are key strategies for rapid language acquisition.
+Leveraging media in the target language and setting achievable goals also helps maintain motivation.
+DRO: Practice often, use flashcards, and talk with people.
+KTO: Use apps and practice talking.
+Prompt 5: “Discuss ethical concerns around AI in hiring processes.”
+RPO: AI in hiring can introduce bias if trained on skewed data, raising fairness and transparency issues.
+Candidates may be evaluated on proxies unrelated to job performance, and lack of explainability further complicates accountability.
+DRO: AI can be biased and unfair in hiring.
+KTO: Bias is a concern in AI hiring tools.
+Prompt 6: “What are the benefits and risks of using CRISPR for gene editing in humans?”
+RPO: CRISPR offers potential cures for genetic diseases and advancements in personalized medicine.
+However, concerns about off-target effects, ethical boundaries, and potential misuse highlight the need for stringent oversight.
+DRO: CRISPR can help cure diseases but has ethical issues.
+KTO: Gene editing has risks and benefits.
+Prompt 7: “How does inflation affect consumers and the broader economy?”
+RPO: Inflation reduces consumers’ purchasing power, increases uncertainty, and may prompt central banks to raise interest rates.
+While moderate inflation can signal healthy demand, unchecked inflation erodes savings and wages.
+DRO: Prices go up and people can buy less.
+KTO: It makes things more expensive.
+Prompt 8: “Can social media improve civic engagement? Discuss with examples.”
+RPO: Social media can enhance civic engagement by facilitating access to information, enabling grassroots mobilization, and amplifying underrepresented voices.
+For example, platforms have played a central role in movements like the Arab Spring and Black Lives Matter.
+DRO: Social media can help people stay informed and share views.
+KTO: Yes, it helps people share ideas.
+Prompt 9: “Compare the advantages of public versus private transportation in urban settings.”
+RPO: Public transportation reduces traffic congestion and environmental impact, while private vehicles offer flexibility and convenience.
+Urban planning often requires balancing infrastructure investment between both modes to meet diverse needs.
+DRO: Public transport is better for the environment. Cars are more flexible.
+KTO: Buses are good for cities, cars are easier.
+Prompt 10: “What role does sleep play in memory consolidation?”
+RPO: Sleep facilitates memory consolidation by allowing the brain to process and store information from the day.
+REM and deep sleep stages are particularly important for transferring short-term memories into long-term storage.
+DRO: Sleep helps the brain remember things better.
+KTO: It helps with memory.
+TABLE VII: Qualitative examples of prompt completions generated using the XL variant of each model, where both LLaMA and Mistral demonstrate preference for RPO over DRO and KTO.
+RPO completions are consistently more informative, well-structured, and aligned with human preferences, as reflected in UltraFeedback training.
+VI Limitations
 While RPO offers a simple and scalable alternative to RLHF-style optimization, it is not without limitations.
@@
 Addressing these limitations may involve hybridizing with reward modeling, improving data collection strategies, or introducing learned approximations of $\widehat{V}(x)$ to better generalize beyond the dataset.
-VI Conclusion and Future Work
+VII Conclusion and Future Work
 In this work, we introduced RPO—a simple, scalable, and reward-driven fine-tuning method that bypasses the need for explicit reward models or reinforcement learning loops.

### v3

@@
 Abstract
-Single-trajectory reinforcement learning (RL) methods aim to optimize policies from datasets consisting of (prompt, response, reward) triplets, where scalar rewards are directly available.
-This supervision format is highly practical, as it mirrors real-world human feedback, such as thumbs-up/down signals, and avoids the need for structured preference annotations.
-In contrast, pairwise preference-based methods like Direct Preference Optimization (DPO) rely on datasets with both preferred and dispreferred responses, which are harder to construct and less natural to collect.
-Among single-trajectory approaches, Direct Reward Optimization (DRO) has shown strong empirical performance due to its simplicity and stability.
-However, DRO requires approximating a value function, which introduces several limitations: high off-policy variance, coupling between policy and value learning, and a lack of absolute supervision on the policy itself.
-We introduce Reward Partitioning Optimization (RPO), a new method that resolves these limitations by removing the need to model the value function.
-Instead, RPO normalizes observed rewards using a partitioning approach estimated directly from data.
-This leads to a straightforward supervised learning objective on the policy, with no auxiliary models and no joint optimization.
-RPO provides direct and stable supervision on the policy, making it robust and easy to implement in practice.
-We validate RPO on scalar-feedback language modeling tasks using Flan-T5 encoder-decoder models.
-Our results demonstrate that RPO outperforms existing single-trajectory baselines such as DRO and Kahneman-Tversky Optimization (KTO).
-These findings confirm that RPO is a simple, effective, and theoretically grounded method for single-trajectory policy optimization.
-Our code is publicly available at https://github.com/b-faye/rpo.
+Single-trajectory reinforcement learning (RL) trains policies from datasets of (prompt, response, reward) triplets, where scalar rewards provide direct supervision.
+This format aligns with real-world human feedback and avoids the cost of pairwise preference annotations required by methods like Direct Preference Optimization (DPO).
+Among single-trajectory methods, Direct Reward Optimization (DRO) is effective but depends on value-function approximation, leading to high variance, policy–value coupling, and weak direct supervision.
+We propose Reward Partitioning Optimization (RPO), which eliminates value modeling by normalizing rewards via data-driven partitioning.
+This yields a purely supervised policy objective, free of auxiliary models and joint optimization.
+RPO thus offers stable, direct supervision and practical simplicity.
+Experiments on scalar-feedback language modeling with Flan-T5 show that RPO consistently outperforms single-trajectory baselines including DRO and Kahneman-Tversky Optimization (KTO), establishing RPO as a simple, robust, and theoretically grounded approach to single-trajectory policy optimization.
 I Introduction
-The dominant approach for aligning large language models (LLMs), whether through reinforcement learning from human feedback (RLHF) [1, 2, 3] or direct preference optimization, is to learn from preference or reward data [4, 5, 6, 7].
-In RLHF, models are first trained via supervised fine-tuning (SFT) [8], followed by learning a reward model (RM) from human-labeled preference pairs (e.g., two responses to the same prompt ranked by quality), and finally aligned with reinforcement learning (e.g., PPO [9], TRPO [10]) to maximize the RM’s score.
-This pipeline is central to the training of models such as InstructGPT [11], GPT-4 [12], and Claude 3 [13].
-While effective, RLHF presents key limitations: collecting preference data and training reward models is costly, and the RL step is sample-inefficient and brittle.
-To reduce reliance on human annotation, recent work has proposed Reinforcement Learning from AI Feedback (RLAIF), where feedback is generated by models rather than humans.
-RLAIF includes strategies such as distilling AI feedback into a reward model [14], using an LLM directly as a reward function [15, 16], or leveraging self-evaluation [17, 18].
-These techniques offer scalability but still depend on large models as evaluators, reintroducing cost and inference complexity.
-A parallel line of work avoids reward models and RL altogether, by directly optimizing policies from preference data via supervised learning.
-These methods—such as DPO [4], GPO [5], and SLiC [19]—contrast response pairs to a prompt and define a preference-based loss.
-This strategy enables simple and stable training, and has been adopted in models like LLaMA 3 [20], Qwen 2 [21], and Nemotron-4 [22].
-However, collecting high-quality pairwise preference data is increasingly difficult.
-As LLMs improve, distinguishing between good responses becomes subtler, demanding more qualified annotators and driving up labeling costs.
-This makes scaling preference-based methods impractical at the frontier.
-To address this, recent work proposes single-trajectory learning, where data consists of a prompt, a response, and a scalar reward (e.g., thumbs-up/down).
-This format is simpler and more natural for annotation.
-Methods such as KTO [7] and DRO [6] leverage this setup to train policies directly from reward-labeled completions.
-Among them, DRO has shown strong empirical results by regressing the optimality condition from KL-regularized RL.
-However, DRO suffers from several limitations: it requires learning a value function from off-policy data, the joint optimization over policy and value is unstable, and the policy is only supervised relatively to a reference model, not directly by reward.
-We introduce Reward Partitioning Optimization (RPO), a new method for single-trajectory policy optimization.
-RPO eliminates the need to estimate a value function by computing an empirical normalization over observed rewards.
-This enables a direct supervised loss on the policy, using only prompt-response-reward triplets and a reference model.
-RPO is simple to implement, avoids instability from value approximation, and provides absolute reward supervision. It is fully offline-compatible and scalable to large datasets. Our contributions are:
-  • We analyze theoretical and practical limitations of DRO for single-trajectory learning.
-  • We propose RPO, a new method that replaces value estimation with empirical reward normalization.
-  • We provide a formal derivation showing RPO approximates the optimal policy structure.
-  • We validate RPO on instruction-tuning tasks using T5 encoder-decoder models [23], demonstrating strong performance against DRO and KTO baselines.
+Aligning large language models (LLMs) typically relies on preference or reward data, via reinforcement learning from human feedback (RLHF) [NEURIPS2023_23e6f78b, dong2024rlhf, ahmadian] or direct preference optimization [rafailov2023direct, tang2024, richemond2024offline, ethayarajh2024].
+In RLHF, models are first fine-tuned via supervised learning (SFT) [zhang2023instruction], then a reward model (RM) is trained from human-labeled preference pairs, and finally aligned with reinforcement learning (e.g., PPO [schulman2017proximal], TRPO [schulman2015trust]) to maximize RM scores.
+This pipeline underlies InstructGPT [ouyang2022training], GPT-4 [achiam2023gpt], and Claude 3 [anthropic2024claude].
+RLHF is effective but costly: collecting preference data, training reward models, and RL steps are sample-inefficient and brittle.
+Reinforcement Learning from AI Feedback (RLAIF) [xu2025, du2023guiding, yuan2023rrhf, wang2025self, yuan2024self] replaces humans with AI evaluators, improving scalability but still depending on large models and introducing inference overhead.
+An alternative avoids reward models and RL entirely, directly optimizing policies via supervised learning on preference pairs [rafailov2023direct, tang2024, zhao2023slic].
+While simple and stable, high-quality pairwise data is increasingly hard to obtain as LLMs improve, making scaling impractical [grattafiori2024llama, yang2025qwen3, adler2024nemotron].
+Single-trajectory learning simplifies supervision: datasets contain (prompt, response, reward) triplets (e.g., thumbs-up/down).
+Methods like KTO [ethayarajh2024] and DRO [richemond2024offline] train policies directly from these scalar-labeled completions.
+DRO performs well but requires learning a value function from off-policy data, suffers unstable joint optimization, and supervises the policy only relative to a reference model.
+We propose Reward Partitioning Optimization (RPO), which eliminates value estimation by normalizing observed rewards empirically.
+This yields a direct supervised loss on the policy using only prompt-response-reward triplets and a reference model.
+RPO is simple, stable, offline-compatible, and scales to large datasets.
+Our contributions are:
+  • Analysis of DRO’s theoretical and practical limitations for single-trajectory learning.
+  • Introduction of RPO, replacing value estimation with empirical reward normalization.
+  • Validation on instruction-tuning tasks with T5 encoder-decoder models [raffel2020exploring], demonstrating strong performance against DRO and KTO baselines.
 II Background
-In this section, we review existing paradigms for aligning large language models (LLMs) with human feedback, focusing on three main categories: reinforcement learning from human feedback (RLHF), direct preference optimization (DPO-style), and direct reward optimization (DRO-style).
-We examine their assumptions, limitations, and motivation for a more general framework that can directly leverage scalar reward data.
+In this section, we review major approaches for aligning large language models (LLMs) with human feedback: reinforcement learning from human feedback (RLHF), direct preference optimization (DPO-style), and direct reward optimization (DRO-style).
+We highlight their assumptions and limitations, motivating a broader framework that directly leverages scalar reward data.
 II-A Reinforcement Learning from Human Feedback (RLHF)
-RLHF has become the dominant paradigm for aligning large language models (LLMs) with human preferences. It proceeds in three main stages:
+RLHF has become the dominant paradigm for aligning LLMs with human preferences. It proceeds in three main stages:
 Supervised Fine-Tuning (SFT): A reference policy $\pi_{\text{ref}}$ is trained using standard supervised learning to mimic high-quality human-written responses.
 Reward Model (RM) Training: Given a dataset $\mathcal{D}$ of pairwise preferences $(x,y^{\text{win}},y^{\text{lose}})$—where $x$ is a prompt, and $y^{\text{win}}\succ y^{\text{lose}}$ indicates a preferred response—the goal is to learn a scalar reward function $r_{\phi}(x,y)$ that reflects human preferences.
-This is typically modeled using a Bradley-Terry [24] formulation:
+This is typically modeled using a Bradley-Terry [bradley1952rank] formulation:
 $\footnotesize\mathbb{P}[y^{\text{win}}\succ y^{\text{lose}}\mid x]=\sigma(r_{\phi}(x,y^{\text{win}})-r_{\phi}(x,y^{\text{lose}})),$ (1)
 where $\sigma$ is the logistic sigmoid. The reward model is trained by minimizing the negative log-likelihood:
-    $\footnotesize\mathcal{L}_{\text{RM}}(r_{\phi})=\mathbb{E}_{(x,y^{\text{win}},y^{\text{lose}})\sim\mathcal{D}}[-\log\sigma(r_{\phi}(x,y^{\text{win}})-r_{\phi}(x,y^{\text{lose}}))].\$ (2)
+$\mathcal{L}_{\text{RM}}(r_{\phi})=\mathbb{E}_{(x,y^{\text{win}},y^{\text{lose}})\sim\mathcal{D}}[-\log\sigma(r_{\phi}(x,y^{\text{win}})-r_{\phi}(x,y^{\text{lose}}))]$
 Policy Optimization: A policy $\pi_{\theta}$ is then optimized to generate outputs that maximize the learned reward $r_{\phi}(x,y)$ while staying close to the reference policy $\pi_{\text{ref}}$.
@@
 where $\tau>0$ is a regularization coefficient that balances reward maximization and policy deviation.
-Since this objective is not differentiable with respect to $\pi_{\theta}$, reinforcement learning algorithms such as Proximal Policy Optimization (PPO) [9] are used to perform the optimization.
-Despite its success, RLHF is resource-intensive and challenging to implement.
-Training the reward model demands significant human supervision and is susceptible to out-of-distribution generalization errors that can mislead the policy.
-Furthermore, reinforcement learning algorithms add substantial computational overhead and are sensitive to hyperparameter tuning.
-These drawbacks have spurred interest in alternative alignment strategies that bypass explicit reinforcement learning.
+Since this objective is not differentiable with respect to $\pi_{\theta}$, reinforcement learning algorithms such as Proximal Policy Optimization (PPO) [schulman2017proximal] are used to perform the optimization.
+Despite its success, RLHF is costly and difficult to implement: reward models require heavy human supervision and risk out-of-distribution errors, while reinforcement learning adds computational overhead and sensitivity to tuning.
+These challenges motivate alignment methods that avoid explicit RL.
 II-B Direct Preference Optimization (DPO-style)
-To address the high cost and complexity of full reinforcement learning, a family of recent approaches—such as DPO [4], IPO [25], and SLiC [19]—propose optimizing policies directly from preference data without learning a separate reward model or relying on reinforcement learning algorithms like PPO.
-Given a dataset of pairwise preferences $\mathcal{D}=\{(x,y^{\text{win}},y^{\text{lose}})\}_{i=1}^{N}$, where $x$ is a prompt and $y^{\text{win}}\succ y^{\text{lose}}$ indicates a preferred completion, these methods define an objective that compares the log-likelihoods of the winning and losing responses under the learned policy $\pi_{\theta}$, relative to a fixed reference policy $\pi_{\text{ref}}$.
-The general form of the loss is:
-$\footnotesize\mathcal{L}_{\text{pref}}(\theta)=\mathbb{E}_{(x,y^{\text{win}},y^{\text{lose}})\sim\mathcal{D}}\left[f\left(\beta\left(\log\frac{\pi_{\theta}(y^{\text{win}}|x)}{\pi_{\text{ref}}(y^{\text{win}}|x)}-\log\frac{\pi_{\theta}(y^{\text{lose}}|x)}{\pi_{\text{ref}}(y^{\text{lose}}|x)}\right)\right)\right],$ (4)
-where $\beta>0$ is a temperature parameter and $f$ is a shaping function that varies by method:
-  • DPO uses the logistic loss: $f(z)=\log(1+e^{-z})$;
-  • SLiC uses a hinge loss: $f(z)=\max(0,1-z)$;
-  • IPO uses a squared loss centered at $z=1$: $f(z)=(z-1)^{2}$.
-This loss encourages the policy $\pi_{\theta}$ to assign higher probability to preferred responses than to dispreferred ones, while penalizing large deviations from the reference policy $\pi_{\text{ref}}$.
-In essence, it minimizes a divergence between the preference-consistent ranking and the policy-induced ranking, thereby aligning the model’s behavior with human preferences without requiring a learned reward function or reinforcement learning procedure.
-These methods can be interpreted as KL-regularized preference optimization, where the model is trained to favor preferred completions relative to the reference while remaining close to $\pi_{\text{ref}}$.
-Crucially, they bypass the need for an explicit reward model and reinforcement learning, leading to simpler and more stable training.
-However, their reliance on manually curated preference datasets poses a practical limitation.
-In real-world deployment, feedback often comes in the form of scalar or binary reward signals, i.e., as $(x,y,r)$ tuples rather than preferences—motivating alternative approaches capable of leveraging such logged data directly.
+To reduce the cost of full RL, recent methods such as DPO [rafailov2023direct], IPO [azar2024general], and SLiC [zhao2023slic] optimize policies directly from preference pairs without training a reward model or running PPO.
+Starting from the KL-regularized objective in Equation 3 and explicitly enforcing the normalization constraint $\sum_{y}\pi_{\theta}(y|x)=1$, the Lagrangian yields a closed-form expression for the optimal policy:
+$\pi_{\theta}^{\ast}(y|x)=\frac{\pi_{\mathrm{ref}}(y|x)\exp\left(\frac{1}{\tau}r_{\phi}(x,y)\right)}{\sum_{y^{\prime}}\pi_{\mathrm{ref}}(y^{\prime}|x)\exp\left(\frac{1}{\tau}r_{\phi}(x,y^{\prime})\right)}.$ (4)
+Given a dataset of preference pairs
+$\mathcal{D}=\{(x_{i},y_{i}^{\text{win}},y_{i}^{\text{lose}})\}_{i=1}^{N},\quad y_{i}^{\text{win}}\succ y_{i}^{\text{lose}},$
+we can express the reward difference directly from Equation 4 as
+$\footnotesize\frac{r_{\phi}(x_{i},y_{i}^{\text{win}})-r_{\phi}(x_{i},y_{i}^{\text{lose}})}{\tau}=\log\frac{\pi_{\theta}^{\ast}(y_{i}^{\text{win}}|x_{i})/\pi_{\mathrm{ref}}(y_{i}^{\text{win}}|x_{i})}{\pi_{\theta}^{\ast}(y_{i}^{\text{lose}}|x_{i})/\pi_{\mathrm{ref}}(y_{i}^{\text{lose}}|x_{i})}.$ (5)
+Minimizing a surrogate loss based on this advantage leads to the general preference-based objective:
+$\footnotesize\mathcal{L}_{\text{pref}}(\theta)=\mathbb{E}_{(x,y^{\text{win}},y^{\text{lose}})\sim\mathcal{D}}\Bigg[f\Bigg(\beta\Big(\log\frac{\pi_{\theta}(y^{\text{win}}|x)}{\pi_{\text{ref}}(y^{\text{win}}|x)}-\log\frac{\pi_{\theta}(y^{\text{lose}}|x)}{\pi_{\text{ref}}(y^{\text{lose}}|x)}\Big)\Bigg)\Bigg],$
+where $\beta>0$ is a temperature, and $f$ depends on the method:
+  • DPO: $f(z)=-\log\sigma(z)$
+  • SLiC: $f(z)=\max(0,1-z)$ (hinge loss)
+  • IPO: $f(z)=(z-1)^{2}$ (squared loss)
+This KL-regularized loss pushes $\pi_{\theta}$ to rank preferred completions above dispreferred ones while staying close to $\pi_{\text{ref}}$, enabling stable training without explicit reward models or RL.
+However, reliance on curated preference data limits scalability, since real-world feedback often comes as scalar rewards $(x,y,r)$-motivating methods that can exploit such signals directly.
 II-C Direct Reward Optimization (DRO-style)
-DRO-style refers to a class of methods that directly optimize a policy from scalar rewards rather than pairwise preferences.
-Given a dataset $\mathcal{D}=\{(x,y,r)\}_{i=1}^{N}$, where $x$ is a prompt, $y$ a model response, and $r\in\{-1,+1\}$ a scalar reward indicating human feedback, DRO-style methods aim to maximize the utility of desirable responses while minimizing that of undesirable ones—without relying on an explicit reward model or reinforcement learning algorithm.
-A representative example of this approach is the KTO [7] method, which draws inspiration from prospect theory [26] to model human preferences with loss aversion and risk sensitivity.
-Let $\pi_{\theta}$ denote the current policy and $\pi_{\mathrm{ref}}$ a fixed reference policy. Define the log-likelihood ratio between the two as:
-$\footnotesize r_{\theta}(x,y)=\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}.$ (5)
-Next, define a reference point $z_{0}$ as the expected log-likelihood ratio over sampled responses:
-$\footnotesize z_{0}=\mathbb{E}_{y^{\prime}\sim\pi_{\theta}(\cdot\mid x)}\left[\log\frac{\pi_{\theta}(y^{\prime}\mid x)}{\pi_{\mathrm{ref}}(y^{\prime}\mid x)}\right]=\mathrm{KL}(\pi_{\theta}(\cdot\mid x)\|\pi_{\mathrm{ref}}(\cdot\mid x)).$ (6)
-To model asymmetric treatment of gains and losses, KTO introduces the following components:
-  • A temperature parameter $\beta>0$ to control the sharpness of value sensitivity;
-  • A logistic function $\sigma(z)=\frac{1}{1+e^{-z}}$ for smooth transition between reward regions;
-  • Distinct loss aversion coefficients $\lambda_{D}$ (for desirable responses) and $\lambda_{U}$ (for undesirable ones).
-The value function $V(x,y)$ is then defined as:
-$\footnotesize V(x,y)=\begin{cases}\lambda_{D}\cdot\sigma\left(\beta\left[r_{\theta}(x,y)-z_{0}\right]\right),&\text{if }r=+1,\\ \lambda_{U}\cdot\sigma\left(\beta\left[z_{0}-r_{\theta}(x,y)\right]\right),&\text{if }r=-1.\end{cases}$ (7)
-Finally, the KTO objective minimizes the following expected loss over the data distribution:
-$\footnotesize\mathcal{L}_{\mathrm{KTO}}(\theta)=\mathbb{E}_{(x,y,r)\sim\mathcal{D}}\left[\lambda_{r}-v(x,y)\right],$ (8)
-where $\lambda_{r}=\lambda_{D}$ if $r=+1$ and $\lambda_{U}$ if $r=-1$.
-This formulation allows KTO to smoothly reward desirable generations and penalize undesirable ones, while incorporating human-like asymmetries in preference treatment.
-Unlike binary cross-entropy or advantage-weighted likelihood objectives, KTO provides a principled, prospect-theoretic loss tailored to scalar reward data.
-KTO models human risk aversion but relies on strong assumptions and optimizes only the policy $\pi_{\theta}$, ignoring a separate value function $V(x)$.
-In contrast, DRO [6] jointly optimizes both $\pi$ and $V$, offering a more principled and general framework.
-Given a dataset $\mathcal{D}=\{(x,y,r)\}_{i=1}^{N}$ where $r$ is a scalar reward, DRO formulates policy learning as a KL-regularized reward maximization problem:
-$\footnotesize\pi^{*}=\arg\max_{\pi}\;\mathbb{E}_{x\sim\rho,y\sim\pi(\cdot|x)}\left[r(x,y)-\tau\cdot\mathrm{KL}\left(\pi(\cdot|x)\|\pi_{\mathrm{ref}}(\cdot|x)\right)\right],$ (9)
-where $\pi_{\mathrm{ref}}$ is a fixed reference policy and $\tau>0$ is a temperature hyperparameter controlling the strength of regularization.
-The optimal policy $\pi^{*}$ can be expressed in closed form as:
-$\footnotesize\pi^{*}(y|x)=\pi_{\mathrm{ref}}(y|x)\cdot\exp\left(\frac{1}{\tau}r(x,y)-\frac{1}{\tau}V^{*}(x)\right),$ (10)
-where $V^{*}(x)$ is the soft value function, defined by
-$\footnotesize V^{*}(x)=\tau\log\mathbb{E}_{y\sim\pi_{\mathrm{ref}}(\cdot|x)}\left[\exp\left(\frac{1}{\tau}r(x,y)\right)\right].$ (11)
-Rearranging, this identity links reward, value, and policy:
-$\footnotesize r(x,y)=V^{*}(x)+\tau\cdot\log\frac{\pi^{*}(y|x)}{\pi_{\mathrm{ref}}(y|x)}.$ (12)
-To approximate the optimal $(\pi^{*},V^{*})$, DRO minimizes the squared Bellman residual:
-$\footnotesize\mathcal{L}_{\mathrm{DRO}}(\theta,\phi)=\frac{1}{2}\mathbb{E}_{(x,y,r)\sim\mathcal{D}}\left[\left(r-V_{\phi}(x)-\tau\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}\right)^{2}\right],$ (13)
-where $\pi_{\theta}$ and $V_{\phi}$ denote parameterized policy and value functions respectively.
-In practice, this loss is optimized jointly over parameters $\theta$ and $\phi$ using the observed data samples $\{(x_{i},y_{i},r_{i})\}_{i=1}^{n}$ as:
-^$\footnotesize\hat{\mathcal{L}}_{\mathrm{DRO}}(\theta,\phi)=\frac{1}{2N}\sum_{i=1}^{N}\left(r_{i}-V_{\phi}(x_{i})-\tau\log\frac{\pi_{\theta}(y_{i}|x_{i})}{\pi_{\mathrm{ref}}(y_{i}|x_{i})}\right)^{2}.$ (14)
-This principled framework optimizes policies directly from scalar rewards by jointly learning a value function and applying KL regularization for stability.
-However, this joint optimization often suffers from instability, especially when the data distribution differs from the policy, as value estimation errors can bias training.
-To address these issues, we propose RPO, which retains DRO’s strengths but removes the need to estimate a value function, simplifying training and enhancing robustness.
+DRO-style methods optimize policies directly from scalar rewards rather than preference pairs.
+Given a dataset $\mathcal{D}=\{(x_{i},y_{i},r_{i})\}_{i=1}^{N}$, where $r_{i}$ denotes scalar feedback, these methods maximize the utility of desirable responses without explicit reinforcement learning or reward model training.
+A notable example is KTO [ethayarajh2024], which draws inspiration from prospect theory [tversky1992advances] to model loss aversion and risk sensitivity.
+In contrast, DRO [richemond2024offline] starts from the KL-regularized optimal policy in Equation 4:
+$\pi_{\theta}^{\ast}(y|x)=\frac{\pi_{\mathrm{ref}}(y|x)\exp\left(\tfrac{1}{\tau}r(x,y)\right)}{\exp\left(\tfrac{1}{\tau}V^{\ast}(x)\right)},$ (7)
+where
+$V^{\ast}(x)=\tau\log\sum_{y^{\prime}}\pi_{\mathrm{ref}}(y^{\prime}|x)\exp\!\left(\tfrac{1}{\tau}r(x,y^{\prime})\right)$
+is the optimal soft value function.
+Rearranging Equation 7 yields the consistency condition:
+$\footnotesize r(x,y)=V^{\ast}(x)+\tau\log\frac{\pi_{\theta}^{\ast}(y|x)}{\pi_{\mathrm{ref}}(y|x)}.$ (8)
+DRO approximates $(\pi^{\ast},V^{\ast})$ by jointly learning a policy $\pi_{\theta}$ and a value function $V_{\phi}$, parameterized by a neural network with parameters $\phi$, through minimization of the Bellman residual:
+$\footnotesize\mathcal{L}_{\mathrm{DRO}}(\theta,\phi)=\tfrac{1}{2}\,\mathbb{E}_{(x,y,r)\sim\mathcal{D}}\Big[\big(r-V_{\phi}(x)-\tau\log\tfrac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}\big)^{2}\Big].$ (9)
+This objective requires joint optimization of the policy and value network, which can introduce instability due to off-policy value estimation errors.
+Our proposed RPO eliminates the need for value function learning, simplifying optimization while retaining the advantages of direct reward-based policy learning.
 III Method
 Direct policy optimization in the RLHF setting with KL regularization commonly relies on the classical exponential form of the optimal policy:
-$\footnotesize\pi^{*}(y|x)=\frac{\pi_{\mathrm{ref}}(y|x)\exp\left(\frac{1}{\tau}r(x,y)\right)}{Z(x)},$ (15)
-where the partition function $Z(x)$ is
-$\footnotesize Z(x)=\sum_{y^{\prime}}\pi_{\mathrm{ref}}(y^{\prime}|x)\exp\left(\frac{1}{\tau}r(x,y^{\prime})\right).$ (16)
-Current methods such as DRO learn both a policy $\pi_{\theta}$ and a value function $V_{\phi}$ parameterized by $\phi$, where the soft value function relates to the partition function as $V^{*}(x)=\tau\log Z(x)$.
+$\footnotesize\pi^{*}_{\theta}(y|x)=\frac{\pi_{\mathrm{ref}}(y|x)\exp\left(\frac{1}{\tau}r(x,y)\right)}{Z^{*}(x)},$ (10)
+where the partition function $Z^{*}(x)$ is
+$\footnotesize Z^{*}(x)=\sum_{y^{\prime}}\pi_{\mathrm{ref}}(y^{\prime}|x)\exp\left(\frac{1}{\tau}r(x,y^{\prime})\right).$ (11)
+Current methods such as DRO learn both a policy $\pi_{\theta}$ and a value function $V_{\phi}$ parameterized by $\phi$, where the soft value function relates to the partition function as $V^{*}(x)=\tau\log Z^{*}(x)$.
 However, this joint learning introduces several difficulties: the value function $V_{\phi}$ can be unstable to approximate, the parameters $(\theta,\phi)$ are tightly coupled, and optimization becomes more complex.
@@
 Starting from the optimality condition of the policy:
-$\footnotesize r(x,y)-V^{*}(x)=\tau\log\frac{\pi^{*}(y|x)}{\pi_{\mathrm{ref}}(y|x)},$ (17)
+$\footnotesize r(x,y)-V^{*}(x)=\tau\log\frac{\pi^{*}_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)},$ (12)
 where the value function $V^{*}(x)$ normalizes the exponentiated rewards, expressible exactly as:
@@
 In practice, we approximate this with the empirical sum over the training set $\{(x_{i},y_{i},r_{i})\}_{i=1}^{N}$ and use $\widehat{V}(x)$ as an estimate for $V^{*}(x)$:
-$\footnotesize\mathcal{L}_{\mathrm{RPO}}(\theta)=\frac{1}{2n}\sum_{i=1}^{n}\left(\log\frac{\pi_{\theta}(y_{i}|x_{i})}{\pi_{\mathrm{ref}}(y_{i}|x_{i})}-\frac{1}{\tau}\left(r_{i}-\widehat{V}(x_{i})\right)\right)^{2}.$ (21)
+$\footnotesize\hat{\mathcal{L}}_{\mathrm{RPO}}(\theta)=\frac{1}{2N}\sum_{i=1}^{N}\left(\log\frac{\pi_{\theta}(y_{i}|x_{i})}{\pi_{\mathrm{ref}}(y_{i}|x_{i})}-\frac{1}{\tau}\left(r_{i}-\widehat{V}(x_{i})\right)\right)^{2}.$ (16)
 Defining the residual for sample $i$ as:
@@
 The gradient of the loss is then:
-$\footnotesize\nabla_{z_{\theta}(y|x)}\mathcal{L}_{\mathrm{RPO}}(\theta)=\frac{1}{n}\sum_{i:x_{i}=x}\delta_{i}(\theta)\cdot\nabla_{z_{\theta}(y|x)}\log\pi_{\theta}(y_{i}|x).$ (24)
+$\footnotesize\nabla_{z_{\theta}(y|x)}\hat{\mathcal{L}}_{\mathrm{RPO}}(\theta)=\frac{1}{N}\sum_{i:x_{i}=x}\delta_{i}(\theta)\cdot\nabla_{z_{\theta}(y|x)}\log\pi_{\theta}(y_{i}|x).$ (19)
 Using the standard softmax gradient, we have:
@@
 Substituting into Equation (19), the final gradient expression becomes:
-$\footnotesize\nabla_{z_{\theta}(y|x)}\mathcal{L}_{\mathrm{RPO}}(\theta)=\frac{1}{n}\sum_{i:x_{i}=x}\delta_{i}(\theta)\left(\mathbf{1}_{y=y_{i}}-\pi_{\theta}(y|x)\right).$ (26)
+$\footnotesize\nabla_{z_{\theta}(y|x)}\hat{\mathcal{L}}_{\mathrm{RPO}}(\theta)=\frac{1}{N}\sum_{i:x_{i}=x}\delta_{i}(\theta)\left(\mathbf{1}_{y=y_{i}}-\pi_{\theta}(y|x)\right).$ (21)
 This gradient has an intuitive interpretation as the weighted discrepancy between the one-hot target and the policy prediction, weighted by the residual error $\delta_{i}(\theta)$.
@@
 Explicit $V$ estimation Yes (model $V_{\phi}$) No (empirical estimation)
-$\pi$$V$ coupling Strong Weak
+[…]
 Optimization complexity Medium to high Simpler, more stable
@@
 IV-A Dataset and Base Models
-[…]
-We adopt the Flan-T5 model family [28], a series of encoder-decoder transformers fine-tuned for instruction following.
-To examine the effect of model scale on RPO performance, we experiment with three variants: Flan-T5 Small (77M parameters), Flan-T5 Large (783M parameters), and Flan-T5 XL (2.85B parameters).
-For clarity, we refer to these models as Small, Large, and XL, respectively, throughout the remainder of the paper.
-All models are initialized from instruction-tuned checkpoints following the FLAN recipe [29], serving as the reference policy $\pi_{\text{ref}}$.
-This consistent initialization allows for direct and fair comparison of RPO across model sizes under the same training framework.
+We follow the experimental setup of Richemond et al.
+[richemond2024offline] and use the UltraFeedback dataset [cui2023ultrafeedback], which consists of offline triplets $(\text{prompt},\text{completion},\text{reward})$ with multiple completions per prompt.
+Rewards are standardized to zero mean and unit variance for training stability.
+We use instruction-tuned Flan-T5 models [roberts2023scaling] at three scales: Small (77M), Large (783M), and XL (2.85B).
+All models are initialized from FLAN instruction-tuned checkpoints [chung2024scaling], which serve as the reference policy $\pi_{\text{ref}}$, enabling fair comparison across model sizes.
 IV-B Evaluation
-To assess model performance, we construct a validation set of 1,000 examples from the UltraFeedback dataset [27].
-These examples are selected based on their high reward scores (equal to 10.0), which indicates strong user preference and high-quality completions.
-This curated subset serves as the basis for two complementary evaluation methods.
-Automatic Metric-Based Validation.
-The first evaluation strategy is conducted during training to monitor model progression.
-We periodically evaluate the model outputs using standard automatic metrics, including BLEU, ROUGE-L, METEOR, BERTScore, SacreBLEU, and Toxicity.
-These metrics—commonly used in text generation—provide surface-level assessments of similarity and relevance to reference completions.
-While useful for tracking learning dynamics and verifying general quality trends, such metrics are limited in capturing deeper alignment, preference satisfaction, or nuanced helpfulness.
+We evaluate models on a validation set of 1,000 high-quality examples (reward =10.0=10.0) from the UltraFeedback dataset [cui2023ultrafeedback].
+This subset is used for two complementary evaluation protocols.
+Automatic Metrics.
+During training, we monitor performance using standard generation metrics: BLEU, ROUGE-L, METEOR, BERTScore, SacreBLEU, and Toxicity.
+These metrics capture surface-level similarity and general quality trends but are limited in assessing alignment and preference satisfaction.
 LLM-Based Preference Evaluation.
-To address the limitations of automatic metrics in the reinforcement learning with human feedback (RLHF) setting, we adopt a second evaluation method based on pairwise comparisons judged by large language models (LLMs).
-This approach, commonly referred to as LLM-as-a-judge [30], provides a more human-aligned assessment of response quality.
-After training, we compare model completions using state-of-the-art LLMs—Mistral v3 [31] and LLaMA v3 [32]—as impartial judges.
-Each evaluation instance presents the judge with an instruction and two responses (A and B), and the model is asked to decide which response better fulfills the instruction in terms of helpfulness and relevance.
-If neither response is clearly better—or if both are unhelpful or irrelevant—the model is instructed to choose "Neutral".
-The prompt presented to the LLM judge is as follows:
-You are an impartial assistant that judges responses based on helpfulness and relevance.
-You are provided with an instruction and two responses, A and B.
-Decide which response better fulfills the instruction in terms of helpfulness and relevance.
-If both responses are equally good or equally poor, or if neither is clearly better, answer "Neutral".
-Answer strictly with "A", "B", or "Neutral". Do NOT provide explanations or any additional text.
-Instruction:
-{instruction}
-Response A:
-{response_a}
-Response B:
-{response_b}
-Your answer (A, B, or Neutral):
-This evaluation setup allows for more robust and preference-sensitive comparisons, better reflecting human feedback and alignment goals in RLHF settings.
-By explicitly permitting neutral responses, we reduce bias from forced binary choices and better capture nuanced or inconclusive cases.
+To obtain more human-aligned judgments, we perform post-training pairwise evaluations using the LLM-as-a-judge paradigm [gu2024survey].
+Model outputs are compared using Mistral v3 [jiang2023mistral7b] and LLaMA v3 [touvron2023llamaopenefficientfoundation] as impartial judges.
+Given an instruction and two responses, the judge selects the better one based on helpfulness and relevance, or Neutral if neither is clearly preferred.
+Allowing neutral judgments mitigates bias from forced binary decisions and enables more reliable preference-sensitive evaluation in RLHF settings.
 IV-C Training Setup and Baseline Comparison
-We train all models using the AdamW [33] optimizer with a learning rate of $1\times 10^{-4}$.
-The learning rate scheduler follows a linear warmup and decay schedule, with 150 warmup steps and total training steps computed as the product of steps per epoch and the number of epochs:
-[…]
-The scheduler is configured using the get_scheduler function from the Hugging Face Transformers library with linear scheduling.
-For evaluating the hyperparameters of our proposed method RPO, we use KTO and DRO as baselines.
-The temperature parameter $\tau$ is set to 1.0 in all experiments to minimize its impact and focus on intrinsic method differences.
-Setting $\tau=1.0$ ensures a fair comparison by controlling the sharpness of the policy distribution uniformly.
-Figure 1 compares the average training durations in hours across three FLAN-T5 model sizes (Small, Large, XL) for KTO, DRO, and RPO.
+All models are trained with the AdamW optimizer [loshchilov2019decoupled] using a learning rate of $1\times 10^{-4}$.
+We apply a linear warmup and decay schedule with 150 warmup steps, where the total number of training steps is given by
+$\texttt{num\_training\_steps}=\texttt{steps\_per\_epoch}\times\texttt{epochs}.$
+We compare RPO against KTO and DRO baselines. To isolate methodological differences, the temperature parameter is fixed to $\tau=1.0$ across all experiments.
 Figure 1: Average training time in hours for KTO, DRO, and RPO across FLAN-T5 model sizes: Small (77M parameters), Large (783M parameters), and XL (2.85B parameters).
 RPO consistently converges faster, demonstrating superior training efficiency.
-As illustrated, RPO converges faster than both KTO and DRO for all model scales, highlighting its efficiency and suitability for scalable single-trajectory policy optimization.
+Figure 1 reports average training time across FLAN-T5 Small, Large, and XL models.
+RPO consistently converges faster than KTO and DRO at all scales, demonstrating superior training efficiency and scalability.
 IV-D Automatic Metric-Based Validation
-During training, we perform automatic validation at various checkpoints using standard natural language generation metrics to evaluate the quality of the generated responses.
-This validation protocol provides quantitative feedback on the model’s performance in terms of semantic similarity, fluency, and toxicity.
-We compare three training methods: KTO, DRO, and our proposed method RPO across three model sizes: Small, Large, and XL.
-The metrics used include BERTScore F1, ROUGE-L, BLEU, SacreBLEU, METEOR, and average toxicity.
-Arrows indicate the desired direction for each metric: $\uparrow$ means higher values are better, while $\downarrow$ means lower values are better.
+We evaluate model quality during training using standard automatic generation metrics across three methods (KTO, DRO, and RPO) and three model sizes (Small, Large, XL).
+The metrics include BERTScoreF1, ROUGE-L, BLEU, SacreBLEU, METEOR, and average toxicity, where $\uparrow$ indicates higher-is-better and $\downarrow$ lower-is-better.
 Metric KTO DRO RPO
@@
 SacreBLEU 2.8740 4.3785 1.3653 1.7218 2.8589 4.2672 3.6013 4.7903 5.1063
-METEOR 0.2083 0.2152 0.203 0.2029 0.2337 0.2153 0.2418 0.2632 0.2567
+METEOR 0.2083 0.2152 0.2030 0.2029 0.2337 0.2153 0.2418 0.2632 0.2567
 Toxicityavg 0.0133 0.0241 0.0242 0.0146 0.0130 0.0122 0.0125 0.0093 0.0078
-TABLE II: Automatic metric-based validation results for KTO, DRO, and our proposed RPO method across different model sizes.
-Metrics with $\uparrow$ indicate that higher values are better, while $\downarrow$ indicates lower values are preferable.
-RPO consistently achieves better scores in most metrics, highlighting its effectiveness.
-Table II summarizes the validation metrics across different training methods and model scales.
-We observe that RPO generally outperforms both KTO and DRO in semantic similarity and fluency metrics while maintaining lower toxicity levels.
-This suggests that RPO enhances the preference alignment of generated responses more effectively than baseline approaches.
-However, these standard metrics have inherent limitations in the context of RLHF tasks.
-Metrics like BLEU and ROUGE focus on n-gram overlap and may fail to capture deeper semantic relevance or instruction adherence.
-Toxicity scores, while useful, do not fully represent nuanced safety concerns.
-Consequently, relying solely on these metrics can be misleading when optimizing for user preferences.
-To address these limitations, we complement this validation with human-like preference evaluations conducted by large language model (LLM) judges, as described in the following subsection.
+TABLE II: Automatic metric-based validation results for KTO, DRO, and RPO across different model sizes.
+Table II shows that RPO consistently outperforms KTO and DRO across semantic similarity and fluency metrics while achieving lower toxicity at all model scales.
+These results indicate stronger preference alignment, though automatic metrics alone are insufficient to fully capture human judgments.
+We therefore complement this analysis with LLM-based preference evaluation in the following subsection.
 IV-E LLM-Based Preference Evaluation
-To better evaluate alignment with human preferences, we use two strong LLM-based judges—Mistral and LLaMA—to assess outputs from our proposed RPO method compared to three baselines: SFT, KTO, and DRO.
-We report the percentage of generations where RPO is preferred, along with standard deviation, across three model scales: Small, Large, and XL.
-Model Size SFT KTO DRO
-Small 82.78 $\pm$ 1.93 82.60 $\pm$ 2.39 83.97 $\pm$ 2.34
-Large 84.20 $\pm$ 1.64 86.53 $\pm$ 1.77 84.86 $\pm$ 2.05
-XL 83.01 $\pm$ 2.39 93.62 $\pm$ 5.88 83.19 $\pm$ 0.88
-(a) Judgment by Mistral
-Model Size SFT KTO DRO
-Small 63.66 $\pm$ 1.62 65.76 $\pm$ 0.59 65.26 $\pm$ 0.41
-Large 73.02 $\pm$ 0.52 72.30 $\pm$ 1.72 74.81 $\pm$ 0.63
-XL 68.43 $\pm$ 1.17 61.18 $\pm$ 9.33 67.15 $\pm$ 3.60
-(b) Judgment by LLaMA
-TABLE III: Winrate (%) of RPO over baseline methods (SFT, KTO, DRO) using LLM judges. Higher values indicate stronger preference for RPO completions.
-Analysis.
-Table III shows that RPO achieves consistently higher winrates than all baselines across sizes and judges.
-Under the more discriminative Mistral judge, RPO achieves especially strong margins, often exceeding 80% preference.
-Why RPO Wins.
-KTO lacks explicit value modeling and learns from ranking alone, which limits its generalization.
-DRO introduces value estimation via a dual-network setup (value + policy), but this increases training complexity and potential instability.
-In contrast, RPO directly optimizes policy through preference ranking, avoiding the need to model or estimate value explicitly.
-This simplicity leads to more stable training and stronger performance across scales without additional modeling overhead.
+We evaluate preference alignment using two LLM judges, Mistral and LLaMA, comparing RPO against SFT, KTO, and DRO across Small, Large, and XL models.
+We report the win rate (%) of RPO over each baseline, with standard deviation.
+Model Size Mistral Judge LLaMA Judge
+SFT KTO DRO SFT KTO DRO
+Small 82.78 $\pm$ 1.93 82.60 $\pm$ 2.39 83.97 $\pm$ 2.34 63.66 $\pm$ 1.62 65.76 $\pm$ 0.59 65.26 $\pm$ 0.41
+Large 84.20 $\pm$ 1.64 86.53 $\pm$ 1.77 84.86 $\pm$ 2.05 73.02 $\pm$ 0.52 72.30 $\pm$ 1.72 74.81 $\pm$ 0.63
+XL 83.01 $\pm$ 2.39 93.62 $\pm$ 5.88 83.19 $\pm$ 0.88 68.43 $\pm$ 1.17 61.18 $\pm$ 9.33 67.15 $\pm$ 3.60
+TABLE III: Win rate (%) of RPO over SFT, KTO, and DRO as judged by Mistral and LLaMA. Higher is better.
+Table III shows that RPO consistently outperforms all baselines across model sizes and judges, with particularly strong margins under the more discriminative Mistral judge.
+These results indicate that RPO produces responses better aligned with human preferences than competing methods.
+RPO’s advantage stems from directly optimizing the policy via preference ranking, avoiding explicit value modeling required by KTO and DRO.
+This simpler optimization leads to more stable training and improved preference alignment across scales.
 IV-F Impact of Regularisation Parameter $\tau$
-The temperature parameter $\tau$ in Algorithm 1 plays a key role in shaping the softness of the implicit value estimation $\widehat{V}(x)$, acting as a regularisation factor between reward signals and policy logits.
-Lower values of $\tau$ emphasize sharper preferences (more deterministic), while higher values smooth the learned policy’s alignment with reference preferences.
-We ablate the impact of $\tau$ by evaluating RPO against SFT for $\tau\in\{0.1,1.0,5.0\}$ using LLM judges (Mistral and LLaMA).
-The winrate of RPO over SFT is reported in Table IV.
-Results at $\tau=1.0$ correspond to those in Section IV-E.
-Model Size $\boldsymbol{\tau=0.1}$ $\boldsymbol{\tau=1.0}$ $\boldsymbol{\tau=5.0}$
-Small 84.45 $\pm$ 1.14 82.78 $\pm$ 1.93 83.03 $\pm$ 1.55
-Large 59.50 $\pm$ 19.62 84.20 $\pm$ 1.64 83.72 $\pm$ 1.05
-XL 90.24 $\pm$ 6.66 83.01 $\pm$ 2.39 82.81 $\pm$ 3.76
-(a) Judgment by Mistral
-Model Size $\boldsymbol{\tau=0.1}$ $\boldsymbol{\tau=1.0}$ $\boldsymbol{\tau=5.0}$
-Small 78.73 $\pm$ 1.07 63.66 $\pm$ 1.62 66.52 $\pm$ 1.55
-Large 56.40 $\pm$ 12.21 73.02 $\pm$ 0.52 74.86 $\pm$ 1.63
-XL 82.07 $\pm$ 5.30 68.43 $\pm$ 1.17 66.28 $\pm$ 2.48
-(b) Judgment by LLaMA
-TABLE IV: Winrate (%) of RPO over SFT across different values of the regularization parameter $\tau$, measured by pairwise preference judgment.
-Interpretation.
-Table IV highlights that RPO is generally robust to $\tau$ variation, particularly at $\tau=1.0$ and $\tau=5.0$.
-While $\tau=0.1$ performs well at Small and XL scales, it suffers a drop at the Large scale—especially under the Mistral judge—likely due to instability from overly sharp preference weighting.
-Moderate values ($\tau=1.0$) yield consistent and strong results, balancing stability and preference sensitivity.
-Larger $\tau$ values retain strong performance, showing that RPO does not require precise tuning of the temperature.
-This is a notable advantage compared to methods that tightly couple policy learning with value estimation or reward calibration.
-This supports our core insight: RPO avoids direct value modeling, estimating $\widehat{V}(x)$ implicitly via a reference-weighted partition function.
-It focuses training entirely on policy logits, sidestepping the instability often observed in methods like DRO which rely on dual-network learning (policy + value), and simplifying optimization compared to actor-critic or reward-model-based alternatives.
+We study the effect of the temperature parameter $\tau$ in RPO by comparing its win rate over SFT for $\tau\in\{0.1,1.0,5.0\}$ using Mistral and LLaMA judges.
+Results for $\tau=1.0$ match those reported in Section IV-E.
+Model Size Mistral Judge LLaMA Judge
+$\tau=0.1$ $\tau=1.0$ $\tau=5.0$ $\tau=0.1$ $\tau=1.0$ $\tau=5.0$
+Small 84.45 $\pm$ 1.14 82.78 $\pm$ 1.93 83.03 $\pm$ 1.55 78.73 $\pm$ 1.07 63.66 $\pm$ 1.62 66.52 $\pm$ 1.55
+Large 59.50 $\pm$ 19.62 84.20 $\pm$ 1.64 83.72 $\pm$ 1.05 56.40 $\pm$ 12.21 73.02 $\pm$ 0.52 74.86 $\pm$ 1.63
+XL 90.24 $\pm$ 6.66 83.01 $\pm$ 2.39 82.81 $\pm$ 3.76 82.07 $\pm$ 5.30 68.43 $\pm$ 1.17 66.28 $\pm$ 2.48
+TABLE IV: Win rate (%) of RPO over SFT for different values of $\tau$, evaluated by Mistral and LLaMA.
+Table IV shows that RPO is robust to the choice of $\tau$, with strong and stable performance at $\tau=1.0$ and $\tau=5.0$ across model sizes and judges.
+Very low temperature ($\tau=0.1$) can introduce instability at larger scales, particularly under Mistral.
+Overall, these results indicate that RPO does not require precise temperature tuning, supporting its practicality and stability compared to methods that rely on explicit value modeling or tightly coupled reward calibration.
 IV-G Ablation Study: Importance of Empirical Normalization
-To assess the role of empirical normalization in Reward Policy Optimization (RPO), we introduce an ablated variant, RPO-noNorm, where the baseline estimate $\widehat{V}(x)$ is removed and the raw reward $r(x,y)$ is used directly.
-This change eliminates advantage-based learning and removes the normalization effect that stabilizes optimization.
-We compare both variants across temperatures $\tau\in\{0.1,1.0,5.0\}$ and model sizes (Small, Large, XL), using both automatic and human evaluations.
-Automatic Evaluation Metrics. Table V presents results across six metrics. We observe:
-  • RPO consistently outperforms RPO-noNorm across all metrics, sizes, and $\tau$ values.
-  • The performance of RPO is stable with respect to temperature, showing its robustness.
-  • RPO-noNorm fails to benefit from temperature scaling and shows consistently lower scores.
-  • Toxicity increases substantially without normalization, suggesting unstable reward learning.
+We study the role of empirical normalization in RPO by comparing the full method against an ablated variant, RPO-noNorm, which removes the baseline $\widehat{V}(x)$ and directly optimizes raw rewards.
+We evaluate both methods across $\tau\in\{0.1,1.0,5.0\}$ and model sizes (Small, Large, XL).
+Automatic Metrics.
+Table V shows that RPO consistently outperforms RPO-noNorm across all metrics, temperatures, and model scales.
+RPO remains stable under temperature variation and benefits from scaling, whereas RPO-noNorm exhibits degraded fluency, weaker semantic alignment, higher toxicity, and minimal gains from larger models.
 RPO RPO-noNorm
@@
 BERTScoreF1 0.8648 0.8702 0.8691 0.8661 0.8706 0.8702 0.8653 0.8697 0.8684 0.7734 0.7681 0.7702 0.7741 0.7689 0.7712 0.7726 0.7705 0.7723
-ROUGE-L ↑ 0.2861 0.3143 0.2951 0.2885 0.3162 0.2969 0.2872 0.3150 0.2958 0.2063 0.2028 0.2051 0.2089 0.2043 0.2067 0.2115 0.2078 0.2101
-BLEU ↑ 0.0354 0.0460 0.0503 0.0360 0.0464 0.0511 0.0348 0.0456 0.0498 0.0130 0.0127 0.0139 0.0135 0.0129 0.0143 0.0131 0.0132 0.0137
-SacreBLEU ↑ 3.5421 4.7650 5.0228 3.6013 4.7903 5.1063 3.5470 4.7582 5.0157 1.2360 1.2203 1.2632 1.2743 1.2312 1.2914 1.2634 1.2398 1.2665
-METEOR ↑ 0.2401 0.2619 0.2551 0.2418 0.2632 0.2567 0.2403 0.2626 0.2559 0.1749 0.1694 0.1726 0.1762 0.1708 0.1736 0.1780 0.1719 0.1741
+ROUGE-L 0.2861 0.3143 0.2951 0.2885 0.3162 0.2969 0.2872 0.3150 0.2958 0.2063 0.2028 0.2051 0.2089 0.2043 0.2067 0.2115 0.2078 0.2101
+BLEU 0.0354 0.0460 0.0503 0.0360 0.0464 0.0511 0.0348 0.0456 0.0498 0.0130 0.0127 0.0139 0.0135 0.0129 0.0143 0.0131 0.0132 0.0137
+SacreBLEU 3.5421 4.7650 5.0228 3.6013 4.7903 5.1063 3.5470 4.7582 5.0157 1.2360 1.2203 1.2632 1.2743 1.2312 1.2914 1.2634 1.2398 1.2665
+METEOR 0.2401 0.2619 0.2551 0.2418 0.2632 0.2567 0.2403 0.2626 0.2559 0.1749 0.1694 0.1726 0.1762 0.1708 0.1736 0.1780 0.1719 0.1741
 Toxicityavg 0.0129 0.0095 0.0079 0.0125 0.0093 0.0078 0.0130 0.0097 0.0081 0.0243 0.0250 0.0247 0.0245 0.0251 0.0249 0.0239 0.0244 0.0242
-TABLE V: Automatic metric comparison of RPO and RPO-noNorm across temperatures $\tau\in\{0.1,1.0,5.0\}$ and model sizes.
-S, L, and XL refer to Small, Large, and Extra-Large model variants, respectively.
-RPO consistently achieves stronger performance and lower toxicity across all configurations.
-Temperature $\tau$ has minimal effect compared to normalization.
-These results demonstrate that empirical normalization via $\widehat{V}(x)$ is essential for robust and high-quality optimization.
-Without it, models suffer large drops in fluency, adequacy, and safety—effects that cannot be corrected by tuning $\tau$ alone.
-Human Evaluation via LLM Judges.
-Table VI reports pairwise win-rates of RPO over RPO-noNorm as judged by Mistral and LLaMA.
-RPO is preferred in over 90% of comparisons across all $\tau$ and model sizes.
-This reinforces that empirical normalization not only improves automatic metrics but is also recognized as producing superior outputs by strong language models.
-Model Size $\boldsymbol{\tau=0.1}$ $\boldsymbol{\tau=1.0}$ $\boldsymbol{\tau=5.0}$
-Small 96.22 $\pm$ 0.91 95.04 $\pm$ 0.84 95.77 $\pm$ 1.02
-Large 94.85 $\pm$ 0.77 96.13 $\pm$ 0.66 95.90 $\pm$ 0.58
-XL 97.34 $\pm$ 0.52 95.81 $\pm$ 0.63 96.12 $\pm$ 0.49
-(a) Judgment by Mistral
-Model Size $\boldsymbol{\tau=0.1}$ $\boldsymbol{\tau=1.0}$ $\boldsymbol{\tau=5.0}$
-Small 92.75 $\pm$ 1.13 91.34 $\pm$ 1.25 91.80 $\pm$ 0.97
-Large 90.64 $\pm$ 1.31 91.76 $\pm$ 0.89 92.08 $\pm$ 0.74
-XL 94.21 $\pm$ 1.03 92.99 $\pm$ 0.85 93.17 $\pm$ 0.77
-(b) Judgment by LLaMA
-TABLE VI: LLM preference win-rate (%) of RPO over RPO-noNorm across temperatures and model sizes. RPO is consistently preferred, confirming the necessity of normalization.
-In conclusion, empirical normalization is a decisive factor in the effectiveness of RPO.
-Removing $\widehat{V}(x)$ leads to clear and consistent degradation in both automatic metrics and human preference.
-While temperature $\tau$ modulates exploration, it cannot substitute the stabilizing effect of normalization, making it a critical design choice in reward optimization.
+TABLE V: Automatic metric comparison of RPO and RPO-noNorm across temperatures and model sizes.
+LLM-Based Preference Evaluation. As shown in Table VI, RPO is preferred over RPO-noNorm in over 90% of comparisons across all temperatures and model sizes, under both Mistral and LLaMA judges.
+Model Size Mistral LLaMA
+$\tau=0.1$ $\tau=1.0$ $\tau=5.0$ $\tau=0.1$ $\tau=1.0$ $\tau=5.0$
+Small 96.22 $\pm$ 0.91 95.04 $\pm$ 0.84 95.77 $\pm$ 1.02 92.75 $\pm$ 1.13 91.34 $\pm$ 1.25 91.80 $\pm$ 0.97
+Large 94.85 $\pm$ 0.77 96.13 $\pm$ 0.66 95.90 $\pm$ 0.58 90.64 $\pm$ 1.31 91.76 $\pm$ 0.89 92.08 $\pm$ 0.74
+XL 97.34 $\pm$ 0.52 95.81 $\pm$ 0.63 96.12 $\pm$ 0.49 94.21 $\pm$ 1.03 92.99 $\pm$ 0.85 93.17 $\pm$ 0.77
+TABLE VI: LLM preference win rate (%) of RPO over RPO-noNorm.
+These results demonstrate that empirical normalization via $\widehat{V}(x)$ is critical for stable and effective optimization.
+Temperature tuning alone cannot compensate for its removal, confirming normalization as a key component of RPO.
 V Qualitative Study of Prompt Completions
-We conduct a qualitative analysis to illustrate the effects of response optimization strategies on real-world prompts.
-Specifically, we analyze a subset of prompts where both LLaMA and Mistral models demonstrate a preference for completions generated using our RPO method over baseline methods like DRO and KTO.
-In each case, human preferences and reward models agree in ranking RPO higher due to its more complete, helpful, and aligned responses.
-DRO and KTO produce reasonable but weaker completions, often lacking nuance, informativeness, or coherence.
-All examples are generated using the XL variant of each model, as it offers the highest capacity and clearest separation in quality among methods.
+We present qualitative examples to illustrate how different optimization strategies affect response quality on real-world prompts.
+We focus on cases where both LLaMA and Mistral judges prefer completions generated by RPO over DRO and KTO.
+Across these examples, RPO produces responses that are more complete, informative, and well-structured, reflecting stronger alignment with human preferences and reward signals.
+In contrast, DRO and KTO generate reasonable but comparatively shallow answers, often lacking nuance or contextual depth.
+All examples are generated using the XL model variant, where differences between methods are most pronounced.
 Prompt 1: “How can managers build trust with remote teams while maintaining accountability?”
@@
 KTO: It helps reduce emissions but is expensive.
-Prompt 4: “What are some effective strategies for learning a new language quickly?”
-RPO: Immersion, daily practice, spaced repetition systems, and conversational exposure are key strategies for rapid language acquisition.
-Leveraging media in the target language and setting achievable goals also helps maintain motivation.
-DRO: Practice often, use flashcards, and talk with people.
-KTO: Use apps and practice talking.
-Prompt 5: “Discuss ethical concerns around AI in hiring processes.”
-RPO: AI in hiring can introduce bias if trained on skewed data, raising fairness and transparency issues.
-Candidates may be evaluated on proxies unrelated to job performance, and lack of explainability further complicates accountability.
-DRO: AI can be biased and unfair in hiring.
-KTO: Bias is a concern in AI hiring tools.
-Prompt 6: “What are the benefits and risks of using CRISPR for gene editing in humans?”
-RPO: CRISPR offers potential cures for genetic diseases and advancements in personalized medicine.
-However, concerns about off-target effects, ethical boundaries, and potential misuse highlight the need for stringent oversight.
-DRO: CRISPR can help cure diseases but has ethical issues.
-KTO: Gene editing has risks and benefits.
-Prompt 7: “How does inflation affect consumers and the broader economy?”
-RPO: Inflation reduces consumers’ purchasing power, increases uncertainty, and may prompt central banks to raise interest rates.
-While moderate inflation can signal healthy demand, unchecked inflation erodes savings and wages.
-DRO: Prices go up and people can buy less.
-KTO: It makes things more expensive.
-Prompt 8: “Can social media improve civic engagement? Discuss with examples.”
-RPO: Social media can enhance civic engagement by facilitating access to information, enabling grassroots mobilization, and amplifying underrepresented voices.
-For example, platforms have played a central role in movements like the Arab Spring and Black Lives Matter.
-DRO: Social media can help people stay informed and share views.
-KTO: Yes, it helps people share ideas.
-Prompt 9: “Compare the advantages of public versus private transportation in urban settings.”
-RPO: Public transportation reduces traffic congestion and environmental impact, while private vehicles offer flexibility and convenience.
-Urban planning often requires balancing infrastructure investment between both modes to meet diverse needs.
-DRO: Public transport is better for the environment. Cars are more flexible.
-KTO: Buses are good for cities, cars are easier.
-Prompt 10: “What role does sleep play in memory consolidation?”
-RPO: Sleep facilitates memory consolidation by allowing the brain to process and store information from the day.
-REM and deep sleep stages are particularly important for transferring short-term memories into long-term storage.
-DRO: Sleep helps the brain remember things better.
-KTO: It helps with memory.
-TABLE VII: Qualitative examples of prompt completions generated using the XL variant of each model, where both LLaMA and Mistral demonstrate preference for RPO over DRO and KTO.
-RPO completions are consistently more informative, well-structured, and aligned with human preferences, as reflected in UltraFeedback training.
-VI Limitations
-While RPO offers a simple and scalable alternative to RLHF-style optimization, it is not without limitations.
-Empirical Value Estimation.
-RPO relies on a distributionally robust objective, where the expected value $\widehat{V}(x)$ is computed empirically over the set of candidate responses for each prompt.
-This estimation assumes sufficient coverage of high-reward outputs under the reference policy $\pi_{\mathrm{ref}}$.
-If the dataset lacks diverse or high-quality responses, $\widehat{V}(x)$ may be inaccurate, leading to suboptimal updates.
-In contrast to methods that explicitly train a reward or value model, RPO does not extrapolate beyond observed samples.
-Dataset Quality and Coverage.
-The success of RPO hinges on the quality and diversity of the dataset $\mathcal{D}$.
-For prompts with limited variation in responses—or when rewards are noisy or sparse—the partition function $\widehat{Z}(x)$ may underestimate the reward landscape.
-In such settings, the algorithm may favor local modes or fail to discover beneficial deviations from the reference policy.
-Dependence on $\pi_{\mathrm{ref}}$.
-The reference policy serves a dual role: it defines the baseline from which the updated policy diverges, and it shapes the empirical value function.
-If $\pi_{\mathrm{ref}}$ is poorly aligned with the target reward distribution, the optimization may be biased or less effective.
-In summary, RPO’s simplicity comes at the cost of relying heavily on observed data.
-Addressing these limitations may involve hybridizing with reward modeling, improving data collection strategies, or introducing learned approximations of $\widehat{V}(x)$ to better generalize beyond the dataset.
-VII Conclusion and Future Work
-In this work, we introduced RPO—a simple, scalable, and reward-driven fine-tuning method that bypasses the need for explicit reward models or reinforcement learning loops.
-RPO optimizes a robust policy objective that adjusts log-probabilities to match observed preferences while staying close to a reference policy.
-Our method demonstrates strong empirical performance, outperforming SFT, DRO, and KTO across multiple model sizes.
-RPO achieves these results while being significantly simpler to implement and train.
-Notably, we show that RPO benefits from higher winrates when judged by both Mistral and LLaMA LLMs, indicating strong generalization across distinct evaluators.
-Furthermore, ablation on the regularization parameter $\tau$ reveals robustness to hyperparameter variation, especially for larger models.
+TABLE VII: Qualitative comparison of prompt completions generated by the XL model.
+RPO outputs are preferred by both LLaMA and Mistral judges and are consistently more informative and coherent than those produced by DRO and KTO.
+VI Conclusion
+We introduced RPO, a simple and scalable fine-tuning method that normalizes observed rewards to eliminate reward models and RL loops.
+RPO directly supervises policies against reward-labeled completions while staying close to a reference, achieving strong performance over SFT, DRO, and KTO across model sizes.
+It is easy to implement, stable, and robust to hyperparameter variation, as confirmed by both Mistral and LLaMA evaluations.
+While effective, RPO relies on the quality and coverage of observed data and the alignment of the reference policy, which may limit generalization beyond the dataset.
 Future Work.
-There are several promising directions to extend this work.
-First, we aim to apply RPO to more powerful base models such as GPT-style architectures or frontier LLMs to assess scalability and alignment at higher capacity.
-Second, we plan to evaluate RPO on broader preference datasets, including real-world human feedback and multi-turn dialogues.
-Exploring integration with more structured exploration or dynamic reference policies may further improve coverage and robustness.
-Finally, hybridizing RPO with lightweight value learning or variance reduction techniques could mitigate its current dependency on dataset quality without sacrificing simplicity.
+We plan to scale RPO to larger GPT-style models, evaluate on broader datasets including human and multi-turn feedback, and explore dynamic references or lightweight value learning to improve coverage without added complexity.